### v1

# OffLight: An Offline Multi-Agent Reinforcement Learning Framework for Traffic Signal Control

Abstract
Efficient traffic signal control (TSC) is essential for modern urban mobility, but traditional systems often struggle to adapt to the complex and dynamic nature of city traffic. While Multi-Agent Reinforcement Learning (MARL) offers promising adaptive solutions, online MARL methods require a significant amount of interactions with the environment, which can be expensive and time consuming. Offline MARL addresses these concerns by leveraging historical traffic data for training, but it faces challenges due to the heterogeneity of behavior policies in real-world datasets—a mix of different controllers makes learning difficult. We introduce OffLight, a novel offline MARL framework specifically designed to handle heterogeneous behavior policies within TSC datasets. OffLight employs Gaussian Mixture Model Variational Graph Autoencoder (GMM-VGAEs) to model the complex distribution of behavior policies, enabling effective learning from diverse data sources. To enhance coordination between agents, we integrate Graph Attention Networks (GATs), allowing agents to make informed decisions based on aggregated information from neighboring intersections. Furthermore, OffLight incorporates Importance Sampling (IS) to correct for differences between the behavior and target policies and utilizes Return-Based Prioritized Sampling (RBPS) to focus on high-quality experiences, thereby improving sample efficiency. Extensive experiments across three real-world urban traffic scenarios—Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections)—demonstrate that OffLight significantly outperforms existing offline RL methods. Notably, OffLight achieves up to a 7.8% reduction in average travel time and an 11.2% decrease in queue length compared to baseline algorithms, particularly in datasets with mixed-quality data. Our ablation studies confirm the effectiveness of OffLight’s components in handling data heterogeneity and enhancing learning performance. These results highlight OffLight’s ability to accurately model heterogeneous behavior policies, mitigate the impact of suboptimal data, and scale to large urban networks. By addressing the critical challenges of offline MARL in TSC, OffLight offers a practical and impactful solution for improving urban traffic management without the risks associated with online learning.
Keywords: Multi-Agent Reinforcement Learning, Offline Reinforcement Learning, Traffic Signal Control, Intelligent Transportation Systems, Deep Reinforcement Learning.
1 Introduction
Efficient traffic signal control (TSC) is important for modern urban mobility, directly affecting congestion levels, travel times, and overall city livability. As urban populations grow and vehicular usage intensifies, traditional traffic management systems struggle to adapt to dynamic and complex traffic flows. Recent advancements in Multi-Agent Reinforcement Learning (MARL) offer promising solutions by enabling decentralized, adaptive, and intelligent control of traffic signals [1, 2].
Figure 1: General Offline MARL Framework for Traffic Signal Control
While online MARL requires agents to interact with the environment in real-time—which can be impractical and unsafe in real-world traffic systems—offline MARL leverages historically collected traffic data to train agents without live experimentation [3]. This approach offers several advantages:
  • Extensive Data Utilization: Utilizes pre-collected traffic data, allowing agents to learn from diverse scenarios and rare events difficult to replicate online [4].
  • Safety and Risk Mitigation: Eliminates risks associated with deploying untested policies in live traffic, ensuring only well-trained agents are implemented [5].
  • Cost Efficiency and Scalability: Reduces the need for expensive simulations or field tests, facilitating deployment across various scales and environments [3].
  • Accelerated Policy Development: Enables rapid iteration and refinement of control policies, expediting research and development [6].
Several offline reinforcement learning (RL) methods have been explored for TSC, often incorporating imitation learning to bootstrap training. For instance, DemoLight uses demonstrations from traditional controllers to initialize policies, enabling efficient exploration and faster convergence compared to pure RL approaches [7]. Cooperative Control for multi-intersection TSC combines imitation learning and deep RL to coordinate multiple intersections, highlighting the benefits of learning from expert demonstrations [8]. CrossLight begins with offline training using behavior cloning and transitions to online RL to adapt policies to new environments, showcasing the effectiveness of imitation learning for strong policy initialization [9]. Building upon these efforts, DataLight introduces an input space modeling technique that captures the spatial distribution of traffic using self-attention mechanisms, demonstrating improved policy performance without requiring live interactions [4]. Figure 1 represents a general framework of an offline RL algorithm.
However, a key challenge in offline MARL for TSC remains unaddressed: real-world datasets often consist of a heterogeneous mix of behavior policies, making it difficult for agents to learn effective strategies. This heterogeneity can arise from a variety of factors, such as:
  • Local Heuristics: Many traffic signals operate on traditional rule-based controllers (e.g., fixed-time or reactive control). These heuristics are often suboptimal and vary between intersections [10].
  • Varying Control Strategies: Over time, traffic systems may switch between different controllers (e.g., from manual control to AI-based systems), resulting in datasets that reflect a blend of expert and suboptimal policies [11].
  • Temporal Variability: Traffic patterns fluctuate throughout the day, leading to variations in the policies being implemented during different time periods. For instance, a traffic signal might switch between morning peak-time strategies and off-peak strategies [10].
This heterogeneity hampers effective learning because:
  • Offline MARL methods are highly sensitive to the underlying distribution of policies in the dataset. Mixing high- and low-quality data complicates the learning process, leading to poor generalization [12].
  • Mixed-policy data can introduce biases, making it harder for the agent to identify the most effective policies [13].
To address these challenges, we introduce OffLight, a novel offline MARL framework that combines Importance Sampling (IS) and Return-Based Prioritized Sampling (RBPS) to mitigate distributional shifts and focus on high-value experiences. This synergy between IS and RBPS enhances both sample efficiency and policy performance.
GMM-VGAE models the diversity in behavior policies within the traffic network, capturing policy heterogeneity by disentangling them in the latent space. By leveraging Graph Neural Networks (GNNs), OffLight combines local observations across intersections into a structured global representation, which is crucial in traffic signal control where agent interactions are highly localized but collectively influence network-wide patterns. This enables OffLight to use IS effectively, correcting for distributional shifts and ensuring stable policy learning from varied data. Meanwhile, RBPS prioritizes high-return episodes, allowing OffLight to focus on the most informative experiences, thus accelerating convergence and optimizing performance.
Key contributions of this paper include:
  • Developing OffLight, an offline MARL framework designed to handle heterogeneous behavior policies in traffic signal control (TSC) using a combination of Importance Sampling (IS) and Return-Based Prioritized Sampling (RBPS).
  • Introducing GMM-VGAE to accurately capture diverse behavior policies, supporting robust learning under policy heterogeneity and enabling effective IS.
  • Leveraging RBPS to prioritize high-return episodes, which enhances sample efficiency and accelerates convergence by focusing on the most informative experiences.
  • Demonstrating OffLight’s scalability and superior performance on real-world traffic datasets, where it consistently outperforms existing methods.
The remainder of the paper is organized as follows. Section 2 reviews related work and provides necessary background on MARL and offline RL in the context of TSC. Section 3 details the OffLight framework, including its architecture and key components. Section 4 describes the experimental setup, and Section 5 presents the results and analysis. Finally, Section 6 concludes the paper and Section LABEL:sec:discussion discusses the key insights gained from this study and limitations of OffLight.
2 Background
2.1 Multi-Agent Reinforcement Learning
Multi-Agent Reinforcement Learning (MARL) involves multiple agents interacting within a shared environment. Each agent learns to maximize the cumulative reward while adapting to the behaviors of other agents. MARL introduces additional challenges such as non-stationarity (due to concurrent learning), the need for coordination, and partial observability, all of which make policy learning significantly harder than in single-agent settings [14, 15].
2.1.1 Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs)
Multi-agent problems are commonly modeled as Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs) [16]. A Dec-POMDP is formally defined by a tuple $\langle N,\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{N},\{\mathcal{O}^{i}\}_{i=1}^{N},P,R,H,\gamma\rangle$, where:
  • $N$ is the number of agents.
  • $\mathcal{S}$ is the set of environment states.
  • $\mathcal{A}^{i}$ is the action space of agent $i$.
  • $\mathcal{O}^{i}$ is the observation space of agent $i$.
  • $P:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is the state transition probability function.
  • $R:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\rightarrow\mathbb{R}$ is the shared reward function.
  • $H$ is the finite horizon.
  • $\gamma\in[0,1)$ is the discount factor.
Each agent $i$ selects actions based on its action-observation history $h^{i}_{t}=\{o^{i}_{0},a^{i}_{0},\dots,o^{i}_{t-1},a^{i}_{t-1},o^{i}_{t}\}$, where $o^{i}_{t}\in\mathcal{O}^{i}$ is the observation at time $t$, and $a^{i}_{t}\in\mathcal{A}^{i}$ is the action. The joint action $\mathbf{a}_{t}=(a_{t}^{1},\dots,a_{t}^{N})$ determines the next state $s_{t+1}$ according to $P$, and all agents receive a shared reward $R$. The objective in Dec-POMDPs is to find a set of decentralized policies $\{\pi^{i}\}_{i=1}^{N}$, where each $\pi^{i}:H^{i}_{t}\rightarrow\mathcal{A}^{i}$ maps action-observation histories to actions, such that the expected cumulative reward is maximized:
$\max_{\{\pi^{i}\}}\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}R(s_{t},\mathbf{a}_{t})\right].$ (1)
2.2 Application to Traffic Signal Control (TSC)
In the context of Traffic Signal Control (TSC), each traffic signal controller can be modeled as an agent within the Decentralized Partially Observable Markov Decision Process (Dec-POMDP) framework [16]. Each agent operates independently and makes decisions based on local observations, such as vehicle counts, vehicle speeds, queue lengths, and current signal phases.
Observation Space: Each traffic signal agent has a limited range of observation. This range reflects the realistic coverage of common sensors such as inductive loop detectors or cameras. For each incoming lane, the agent observes:
  • The number of vehicles $\{n_{l}\}_{l=1}^{L_{i}}$, where $L_{i}$ is the number of lanes for traffic signal $i$.
  • The average speed of vehicles $\{s_{l}\}_{l=1}^{L_{i}}$, normalized by the speed limit.
  • The number of halted vehicles, representing the queue length $\{q_{l}\}_{l=1}^{L_{i}}$.
  • The current phase ID, indicating the traffic signal’s current phase.
These local observations provide the sensory input required for each agent to make real-time decisions about the traffic signal phases.
Action Space: Each traffic signal agent controls the signal phases at its intersection. The available actions correspond to selecting one of the predefined green phases from the list of possible phases for that intersection. An agent may select a green phase and it must then transition to the yellow phase, which is enforced by the environment. The action selection interval is fixed at 5 simulation seconds.
Reward Function: A commonly used reward metric is queue length. Minimizing the queue length across all lanes in the network directly reflects the reduction of vehicle stoppage. The global reward at each time step $r_{t}$ is defined as:
$r_{t}=-\sum_{l\in L}q_{l}(t),$
where $q_{l}(t)$ represents the queue length on lane $l$ at time $t$, and $L$ is the set of all lanes in the network. This reward structure encourages agents to minimize congestion and improve traffic flow throughout the system.
Deep MARL: In multi-agent TSC, the actions of one traffic signal directly impact the observations and performance of neighboring intersections. Therefore, coordination between agents is crucial to achieve optimal traffic management [17]. Traditional approaches, such as Independent Q-Learning (IQL) [18], treat each agent as independent, which may lead to suboptimal coordination. Recent advancements in the application of deep MARL in TSC include frameworks such as CoLight [11], PressLight [19], AttendLight [20] which utilize Graph Attention Networks (GATs) [21, 11] to model inter-agent communication. These frameworks model the traffic network as a graph, where intersections are treated as nodes and roads as edges, allowing traffic signals to share information with neighboring intersections. By focusing attention on the most relevant neighbors, these methods have achieved significant success in optimizing traffic flow through coordinated agent behavior.
Following the success of these frameworks in the online MARL setting, we choose this approach as the foundation for our framework to model offline MARL policies. By leveraging the well-established GAT-based communication mechanism, we aim to maintain effective coordination between agents, even when learning occurs from static offline datasets. This approach ensures that OffLight can efficiently handle the complexity of real-world traffic systems, while addressing the challenges inherent in offline learning.
2.3 Offline Multi-Agent Reinforcement Learning
Offline Multi-Agent Reinforcement Learning involves learning effective policies from a fixed dataset of interactions without further environmental exploration during training [3, 4]. This setting is crucial in scenarios where data collection is expensive, risky, or impractical, such as traffic signal control. Offline MARL introduces unique challenges:
  • Distributional Shift: The static dataset may not adequately represent the state-action distribution encountered when executing the learned policies, leading to overestimation of out-of-distribution actions and degraded performance [22].
  • Extrapolation Error: Without any interaction with the environment, agents must generalize based solely on the dataset provided. This limitation risks extrapolation errors, where agents might attempt actions unsupported by the data, leading to poor or unsafe decisions when deployed.
  • Limited Coverage of Optimal Strategies: Offline datasets may lack the comprehensive coverage needed for optimal joint policies. This is particularly challenging in MARL, as optimal coordination patterns often require data from multiple interacting policies, which may be missing in historical data.
Recent offline RL algorithms, such as Conservative Q-Learning (CQL) [12] and TD3+BC [23], have been developed to address these challenges by regularizing learned policies and penalizing overestimation of Q-values for unseen actions. While originally designed for single-agent settings, these methods can be extended to multi-agent systems using frameworks like Centralized Training with Decentralized Execution (CTDE) [24].
2.3.1 Conservative Q-Learning (CQL)
Conservative Q-Learning (CQL) [12] is an offline RL algorithm designed to address the overestimation of out-of-distribution (OOD) actions by learning conservative Q-value estimates. CQL modifies the standard Bellman error minimization by adding a regularization term that penalizes Q-values for actions not present in the dataset. The CQL objective is: {dmath} min_Q α( E_s ∼D, a ∼π(a|s) [Q(s,a)] - E_(s,a) ∼D [Q(s,a)] ) + L_Bellman(Q), where $\mathcal{D}$ is the offline dataset, $\pi(a|s)$ is a policy (e.g., the uniform policy), $\alpha$ is a regularization coefficient, and $\mathcal{L}_{\text{Bellman}}(Q)$ is the standard Bellman error. In multi-agent settings, the joint action space grows exponentially with the number of agents, making it infeasible to represent the full joint Q-function. To address this, CQL can be extended using factorization methods or the CTDE paradigm [24]. CTDE allows agents to be trained centrally with access to global information, while at execution time, they act based on local observations, enabling scalability and coordination among agents even with limited data.
2.3.2 TD3+BC
TD3+BC [23] enhances the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm [25] by integrating a behavior cloning (BC) term into the policy update. This addition encourages the learned policy to stay close to the behavior policy present in the offline dataset, mitigating issues arising from distributional shift. The modified policy objective in TD3+BC is:
$\min_{\pi_{\theta}}-\mathbb{E}_{s\sim\mathcal{D}}[Q_{\phi}(s,\pi_{\theta}(s))]+\lambda\mathbb{E}_{(s,a)\sim\mathcal{D}}[\|\pi_{\theta}(s)-a\|_{2}^{2}],$ (2)
where $\pi_{\theta}$ is the policy network parameterized by $\theta$, $Q_{\phi}$ is the critic network parameterized by $\phi$, and $\lambda$ balances the importance of the behavior cloning term. In multi-agent systems, TD3+BC can be adapted by allowing each agent to learn its own policy while considering the interactions with other agents. Using the CTDE framework [26], agents are trained with access to additional global information (e.g., other agents’ actions or policies) to improve coordination. During execution, agents operate based on their local observations, ensuring scalability and feasibility in decentralized environments.
2.4 Challenge of Heterogeneous Behavior Policies in Offline MARL
A unique challenge in offline MARL, particularly in domains like Traffic Signal Control (TSC), is managing heterogeneous behavior policies present in the offline dataset. In TSC, data often come from various sources, times, or operational contexts, leading to a mix of behavior policies that agents must learn from. This heterogeneity adds complexity to offline MARL in the following ways:
  • Increased Risk of Distributional Shift: When policies in the offline data vary significantly, the likelihood of a distributional mismatch between the training data and the learned policy increases. The agents are more prone to encountering out-of-distribution actions, resulting in unreliable Q-value estimates and degraded policy performance, as the diversity in behavior data can make distributional correction methods less effective [22].
  • Difficulty in Behavior Policy Estimation: For methods relying on importance sampling or policy constraints, accurately estimating a single behavior policy becomes challenging when the data reflects multiple, potentially conflicting policies. Modeling this mixture accurately is essential to effective learning but is significantly more complex in offline MARL with heterogeneous datasets [13].
Addressing the challenge of heterogeneous behavior policies is essential for effective offline MARL, especially in domains like TSC where operational conditions and policies can vary widely.
2.5 Motivation for the OffLight Framework
To tackle the specific challenge of heterogeneous behavior policies in offline MARL, we propose the OffLight framework. OffLight is designed to:
  • Model Heterogeneous Behavior Policies: Utilize advanced modeling techniques, such as Variational Graph Auto-Encoders (GMM-VGAEs), to accurately estimate and represent the diverse behavior policies present in the offline dataset.
  • Mitigate Distributional Shift: Employ importance sampling with the estimated behavior policy to correct for discrepancies between the behavior policies and the target policy, reducing overestimation and improving policy evaluation.
  • Improve Sample Efficiency: In addition to IS and GMM-VGAE, OffLight integrates Return-Based Prioritized Sampling (RBPS) based on episodic returns. RBPS prioritizes high-reward episodes, ensuring that the learning process focuses on the most informative and successful experiences. This prioritization accelerates convergence and enhances the robustness of the learned policies.
By focusing on accurately modeling and leveraging the heterogeneous behavior policies in the offline data, OffLight addresses a critical challenge in offline MARL for TSC.
3 OffLight Framework
OffLight is designed to tackle the challenges of offline multi-agent reinforcement learning (MARL) in the domain of traffic signal control (TSC). It combines a self-supervised learning approach with importance sampling (IS) to account for heterogeneous policies and improve learning from static, pre-collected datasets. IS and RBPS weights are collected before the offline RL training and stored. During the offline RL training phase, these weights are used to improve learning. Below, we outline the architecture and key components of OffLight, highlighting how it handles mixed-policy data to enhance traffic management performance.
Figure 2: OffLight Architecture: Integrating GMM-VGAE with IS and RBPS for offline MARL in traffic signal control
3.1 Dataset Structure
We consider a dataset $\mathcal{D}$ collected from a network of $N$ intersections over $M$ episodic trajectories. Each trajectory $\tau_{k}$ comprises a sequence of interactions between the agents (traffic signals) and the environment. Formally, the dataset is represented as:
$\mathcal{D}=\left\{\tau_{k}\right\}_{k=1}^{M},\quad\text{where}\quad\tau_{i}=\left\{\left(\mathcal{O}_{t},\mathcal{A}_{t},\mathcal{O}^{\prime}_{t},\mathcal{R}_{t}\right)\right\}_{t=0}^{T_{k}},$ (3)
where:
  • $\mathcal{O}_{t}=\left\{o_{t}^{1},o_{t}^{2},\dots,o_{t}^{N}\right\}$ denotes the observations at time step $t$ for each of the $N$ intersections.
  • $\mathcal{A}_{t}=\left\{a_{t}^{1},a_{t}^{2},\dots,a_{t}^{N}\right\}$ represents the joint actions taken by the agents at time $t$.
  • $\mathcal{O}^{\prime}_{t}=\left\{o_{t+1}^{1},o_{t+1}^{2},\dots,o_{t+1}^{N}\right\}$ are the subsequent observations after taking actions $\mathcal{A}_{t}$.
  • $\mathcal{R}_{t}$ is the reward received at time $t$.
  • $T_{k}$ is the length of trajectory $\tau_{k}$.
This structure reflects real-world data collected from sensor-equipped intersections and enables learning traffic control policies from historical data.
3.2 Self-Supervised Learning of Behavior Policies
To effectively model diverse behavior policies in offline traffic signal control (TSC) data, OffLight uses a Gaussian Mixture Model-based Variational Graph Autoencoder (GMM-VGAE), designed to capture both spatial and temporal dependencies while addressing policy heterogeneity across intersections and time.
  • GMM for Policy Heterogeneity: The GMM structure in the VGAE models the latent space as a mixture of Gaussian distributions, allowing OffLight to represent multiple underlying policies. This disentanglement of varied traffic control behaviors within the latent space enables OffLight to more accurately capture and leverage diverse policies from real-world datasets.
  • Graph Attention Networks (GATs) and LSTM for Spatial-Temporal Dynamics: GATs capture spatial dependencies by allowing each agent to attend to relevant neighboring intersections. This attention mechanism is structured to combine local information effectively, creating a global latent representation of behavior policies across the network. Additionally, GATs enable efficient parameter sharing across agents, improving scalability for TSC in a manner similar to MARL by processing local data in a structured and computationally manageable way. LSTM layers capture the temporal evolution of traffic patterns by processing sequences of observations over time, preserving the dynamic dependencies needed for effective policy modeling.
Together, these components enable GMM-VGAE to accurately model both local and global interactions in the traffic network, ensuring OffLight’s robustness to the diverse and dynamic nature of traffic conditions. This architecture makes OffLight highly scalable and effective in learning from heterogeneous TSC datasets.
By constructing a latent space with Gaussian mixtures and capturing spatial-temporal dependencies through GAT and LSTM layers, OffLight provides a structured way to estimate behavior policies at each intersection. This enables seamless integration of importance sampling (IS) and Return-Based Prioritized Sampling (RBPS). Together, these techniques enhance learning from offline datasets by prioritizing transitions aligned with the target policy and derived from high-reward episodes. Below, we discuss how IS and RBPS are applied within OffLight to balance distributional shifts and improve sample efficiency, advancing offline reinforcement learning for traffic signal control (TSC).
Encoder: Captures spatial dependencies using Graph Attention Networks (GATs) and temporal dynamics with a Long Short Term Memory (LSTM) layer. This allows the model to effectively process both the local and global traffic behavior.
$\mathbf{z}=\text{Encoder}(\mathcal{O}_{t},\mathcal{A}_{t-1},\mathbf{A})$
Latent Space: Represents the underlying policies as a mixture of Gaussian variables, enabling the disentanglement of different policy behaviors across the network.
$p(\mathbf{z})=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{z}\mid\mu_{k},\Sigma_{k})$
Decoder: Reconstructs the policy distribution based on the latent representation and current observation, providing an estimate of the behavior policy for each traffic signal.
$\mathcal{A}_{t}=\text{Decoder}(\mathcal{O}_{t},\mathbf{z}_{t},\mathbf{A})$
represents the estimated probability distribution $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\cdot)$ over actions for agent $n$ at time step $t$, conditioned on the current observation and global latent policy representation.
A benefit of using GATs is that parameters can be shared across agents, thereby reducing model complexity and improving the learning efficiency. By learning these latent representations, OffLight can accurately model heterogeneous policies and ensure that the learning process accounts for diverse traffic control strategies. A detailed explanation of GMM-VGAE can be found in Appendix A.3.
3.3 Importance sampling Integration
To address the challenge of distributional shift between the behavior policy and the target policy, OffLight integrates importance sampling (IS). This mechanism adjusts the influence of each transition based on its alignment with the target policy, ensuring that the learning algorithm emphasizes relevant and high-quality data.
$w_{\text{IS},t}^{k}=\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (4)
where:
  • $w_{\text{IS},t}^{k}$ is the importance sampling weight for the transition at time step $t$ in episode $k$.
[…]
  • $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})$ represents the estimated behavior policy probability for agent $n$, obtained from the GMM-VGAE, conditioned on it’s hidden state $\mathbf{h}_{t}^{i}$ and the latent embedding of behavior policy $\mathbf{z}_{t}$.
Importance sampling corrects for the discrepancy between the behavior policy and the target policy by re-weighting transitions  [^2] The target policy is the policy that the offline RL algorithm seeks to optimize, while the behavior policy refers to the policy used to generate the offline dataset.. Transitions where the actions are more probable under the target policy relative to the behavior policy receive higher importance sampling weights.
3.4 Return-Based Prioritized Sampling
To enhance sample efficiency and accelerate the learning process, OffLight employs Return-Based Prioritized Sampling (RBPS), similar to [27]. This strategy prioritizes episodes based on their cumulative rewards, ensuring that the learning algorithm focuses on more successful traffic control experiences.
$w_{\text{RBPS}}^{k}=C\left(\frac{G_{k}-G_{\text{min}}}{G_{\text{max}}-G_{\text{min}}}+p_{\text{base}}\right)$ (5)
where:
  • $w_{\text{RBPS}}^{k}$ is the Return-Based Prioritized Sampling weight for episode $k$.
  • $G_{k}=\sum_{t=0}^{T_{k}}r_{t}$ is the total reward for episode $k$.
  • $G_{\text{min}}$ and $G_{\text{max}}$ are the minimum and maximum total rewards across all episodes in the dataset, respectively.
  • $p_{\text{base}}$ is a small positive constant added to ensure that all episodes have a non-zero probability of being sampled.
  • $C$ is a normalization constant to maintain numerical stability or ensure the weights sum to a desired value.
RBPS ensures that episodes with higher returns, indicative of more effective traffic management, are sampled more frequently. This focus accelerates learning by emphasizing experiences that contribute significantly to improved traffic flow and reduced congestion. This especially proves beneficial when the distribution of the data is skewed or multimodal as can be seen in Figure 3.
(a) Jinan
(b) Hangzhou
(c) Manhattan
Figure 3: Distribution of Episodic Returns in mixed-policy datasets. It highlights the variability and heterogeneity in the offline dataset, with some episodes achieving high returns while others are suboptimal. RBPS targets this imbalance by prioritizing episodes with higher returns, ensuring that learning is focused on successful traffic control strategies.
3.5 Combining importance sampling and Return-Based Prioritized Sampling
Figure 4: Traffic networks used in the experiments: (a) Jinan, (b) Hangzhou, and (c) Manhattan, illustrating the varying scales and complexities of the test scenarios.
OffLight synergistically combines importance sampling (IS) and Return-Based Prioritized Sampling (RBPS) to address both distributional shift and sample efficiency in offline MARL for TSC. This combined weighting scheme ensures that the learning algorithm emphasizes transitions that are both aligned with the target policy and derived from high-reward episodes, thereby enhancing the effectiveness and robustness of policy learning.
$w_{\text{combined},t}^{k}=w_{\text{IS},t}^{k}\times w_{\text{RBPS}}^{k}$ (6)
importance sampling adjusts the weights based on policy alignment, while RBPS prioritizes transitions from successful episodes, ensuring that the model learns from both relevant and high-quality data.  [^3] Multiplying $w_{\text{IS},t}^{k}$ and $w_{\text{RBPS}}^{k}$ ensures that only transitions that are both relevant to the target policy and come from successful episodes are given significant weight. This dual prioritization enhances the effectiveness and robustness of policy learning.
Normalization: After combining, the weights are normalized within each minibatch to maintain their probabilistic interpretation and ensure numerical stability.
$\tilde{w}_{\text{combined},t}^{k}=\frac{w_{\text{combined},t}^{k}}{\sum_{j}w_{\text{combined},t}^{j}}$ (7)
Clipping: To prevent extremely large weights from destabilizing the training process, the combined weights are clamped to a maximum value.
[…]
Integration into Loss Functions: The normalized and clamped combined weights are applied to the Temporal Difference (TD) error and the Conservative Q-Learning (CQL) loss to adjust the influence of each transition during training.
3.6 Integration with Offline RL Algorithms
OffLight can easily be integrated with existing offline RL algorithms. CQL and TD3+BC are chosen as baseline offline RL algorithms. The combined weighting scheme—comprising importance sampling (IS) weights and Return-Based Prioritized Sampling (RBPS) weights—is incorporated into both the loss functions and the sampling mechanisms of these algorithms to ensure effective bias correction and sample prioritization.
Loss Function Adjustment: The combined weights $\tilde{w}_{\text{combined},t}$ are used to scale the loss components, ensuring that transitions from high-reward episodes aligned with the target policy have a greater impact on policy and value updates. This scaling is formalized in the loss functions of CQL and TD3+BC as follows:
$\mathcal{L}=\mathbb{E}_{(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\sim\mathcal{D}}\left[\tilde{w}_{\text{combined},t}\cdot\ell(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\right],$ (9)
where $\ell(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})$ represents the specific loss term associated with the chosen RL algorithm (e.g., Bellman error for CQL or policy loss for TD3+BC).
Sampling Mechanism: Episodes are sampled based on their total rewards, prioritizing those with higher cumulative returns. This is implemented by adjusting the sampling probabilities $p_{i}$ according to $w_{\text{RBPS}}$, ensuring that more successful episodes are sampled more frequently. This prioritization enhances the efficiency of the learning process by focusing on the most informative and effective experiences.
By integrating these techniques, OffLight provides a robust framework for learning effective traffic control policies from static datasets, even in the presence of policy heterogeneity.
4 Experimental Setup
We conduct experiments to evaluate OffLight’s performance across three real-world urban traffic scenarios: Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections). These scenarios vary in complexity and traffic demand, allowing for a comprehensive evaluation of OffLight’s scalability and robustness.
The experiments are designed to address the following key questions:
  1. 1.
    How effectively does OffLight model heterogeneous behavior policies in offline datasets?
  2. 2.
    What is the impact of incorporating importance sampling on learning from mixed-policy datasets?
  3. 3.
    How does the proportion of suboptimal policy data affect overall performance?
4.1 Traffic Scenarios
Average Network
Demand (veh/hr)
Low Medium High Intersections
Jinan 450 550 650 12
Hangzhou 350 420 500 16
Manhattan 150 250 350 196
Table 1: Traffic demand levels and network size for the three urban traffic scenarios used in the experiments, including Jinan, Hangzhou, and Manhattan.
The experiments include three distinct traffic networks: Jinan, a small network with moderate traffic complexity (12 intersections); Hangzhou, a medium network with higher traffic flow (16 intersections); and Manhattan, a large urban network (196 intersections) with complex traffic patterns. Each scenario captures different levels of complexity, simulating low, medium, and high traffic demand conditions as shown in Table 1.
4.2 Data Collection
The dataset used for training includes traffic scenarios generated using a variety of controllers, including traditional rule-based methods (Fixed Time, Greedy, Max Pressure, Self-Organizing Traffic Light (SOTL)) and expert reinforcement learning controllers (MAPPO). Random policy controllers are also used to simulate suboptimal behavior. Each simulation runs for 1 hour, divided into 10 episodes of 6 minutes each, across three traffic demand levels (low, medium, and high). This results in 100 hours of simulation per controller for each demand level.
4.3 Baselines
OffLight’s performance is compared against several baseline offline reinforcement learning methods: Behavior Cloning (BC), Conservative Q-Learning (CQL), and TD3+BC. Each algorithm is trained for $20$k timesteps.
4.4 Evaluation Metrics
Performance is evaluated using two key metrics: Average Travel Time (ATT) and Queue Length (QL). These metrics are calculated across different traffic demand levels to assess the effectiveness of OffLight in managing traffic flow and reducing congestion.
5 Results and Discussion
This section presents the evaluation of OffLight across various traffic scenarios. We measure performance in terms of Average Travel Time (ATT) and Queue Length (QL), highlighting the effectiveness of OffLight in learning from mixed-policy datasets.
5.1 Performance on Mixed-Quality Data
We assess OffLight’s performance on datasets with varying proportions of expert and random policy data. The results are shown for three traffic networks: Jinan, Hangzhou, and Manhattan, under low, medium, and high traffic demand.
Category Algorithm Jinan Hangzhou Manhattan
Low Medium High Low Medium High Low Medium High
Rule-based Random 411.97 450.61 470.82 418.89 430.01 451.15 774.33 784.56 790.14
Fixed Time 354.95 380.26 407.25 370.11 392.69 407.33 539.07 574.17 612.77
Greedy 287.10 310.08 337.60 303.02 307.41 314.72 484.01 527.01 575.03
Max Pressure 386.43 409.47 434.91 405.95 427.04 437.44 640.96 658.31 684.59
SOTL 399.30 418.34 443.35 447.22 470.14 476.91 540.55 583.13 615.19
Online MAPPO 263.87 282.86 301.49 272.57 280.82 292.72 322.99 367.83 402.99
Offline BC 364.50 388.18 409.36 382.59 393.38 408.46 561.07 595.22 617.59
TD3+BC 321.87 356.51 364.26 324.18 338.52 364.97 489.78 495.13 545.11
CQL 326.06 337.74 372.41 369.63 350.06 367.33 495.29 524.25 582.87
OffLight (TD3+BC) 292.90 317.29 316.91 288.52 299.59 313.87 440.80 423.34 441.54
OffLight (CQL) 303.24 307.34 324.00 325.27 304.55 308.56 439.82 456.10 483.78
Table 2: Average Travel Time (ATT): Mixed Policy Dataset
Category Algorithm Jinan Hangzhou Manhattan
Low Medium High Low Medium High Low Medium High
Rule-based Random 590.14 809.45 944.45 260.09 295.46 357.30 2899.23 3291.49 3590.03
Fixed Time 411.88 553.15 716.47 193.45 231.09 270.04 615.94 757.01 884.03
Greedy 216.56 356.59 514.37 170.81 192.55 219.85 512.27 676.02 830.96
Max Pressure 500.51 639.20 803.91 254.08 309.32 343.49 758.94 881.23 991.15
SOTL 518.57 645.27 790.08 320.91 377.54 418.35 607.80 767.82 886.84
Online MAPPO 170.98 213.19 228.95 152.81 162.34 171.91 374.44 452.74 512.33
Offline BC 417.35 554.59 689.67 233.26 269.77 308.28 971.47 1173.25 1303.26
TD3+BC 372.31 509.34 586.35 193.78 240.24 273.07 874.53 956.15 1129.03
CQL 369.72 477.32 566.41 227.52 240.24 275.83 864.52 1023.94 1218.43
OffLight (TD3+BC) 312.88 423.88 493.12 169.07 222.17 222.61 806.83 944.64 962.81
OffLight (CQL) 314.86 402.17 553.09 189.50 219.46 240.54 798.09 979.13 999.89
Table 3: Queue Length (QL): Mixed Policy Dataset
Figure 5: Performance Comparison of OffLight (CQL) and OffLight (TD3+BC) on Mixed Data showing the improvements in average travel time (ATT) across different traffic demand levels
5.1.1 Average Travel Time (ATT)
Figure 5 compares the performance of OffLight with baseline offline RL methods across three traffic scenarios—Jinan, Hangzhou, and Manhattan—at low, medium, and high traffic demand levels. Lower ATT values indicate more efficient traffic flow.
Across all three scenarios (Jinan, Hangzhou, Manhattan), OffLight consistently outperforms both TD3+BC and CQL baselines. In particular, OffLight (TD3+BC) shows the best performance under high traffic demand conditions, particularly in the Jinan and Manhattan scenarios. In the Manhattan scenario, where traffic complexity is the highest, OffLight achieved up to a 6.9% reduction in ATT at high demand levels when integrated with CQL, and a 7.8% reduction when integrated with TD3+BC, as compared to their respective baselines. This highlights OffLight’s ability to handle large-scale networks effectively under congested conditions. In low traffic demand scenarios, the performance improvements were less pronounced, which could indicate that under lighter traffic, the margin for improvement is naturally smaller. The results suggest that OffLight’s incorporation of IS and RBPS allows it to prioritize transitions that are most relevant to traffic efficiency, even in datasets containing a mix of expert and suboptimal policies. By focusing learning on higher-quality data, OffLight is able to reduce congestion more effectively, particularly in high-demand, complex networks like Manhattan.
5.1.2 Queue Length (QL)
Figure 6: Learning curves for the Jinan scenario under medium traffic demand, showing the convergence of queue length for OffLight and baseline methods.
Table 2 presents the queue length results for the three traffic scenarios, indicating the efficiency of OffLight in reducing the number of vehicles queued at intersections. OffLight (TD3+BC) shows significant reductions in queue lengths, particularly in medium to high traffic demand scenarios. For example, in the Hangzhou scenario with medium traffic, OffLight (TD3+BC) reduced the queue length by 11.2% compared to TD3+BC alone. In Manhattan, OffLight also showed marked reductions in queue length, outperforming both baselines across all traffic levels. For instance, OffLight (CQL) achieved a 6% reduction in queue length under high traffic demand, which is critical in reducing congestion in highly complex urban networks. However, under low traffic demand, queue length reductions are less pronounced, likely because fewer vehicles are on the road, making it easier for any control policy to manage traffic effectively. The strong performance in queue length reduction, particularly in medium and high traffic scenarios, highlights OffLight’s ability to coordinate intersections more effectively. The IS mechanism ensures that OffLight learns to prioritize actions that lead to better overall traffic management, minimizing stoppages at intersections.
Figure 6 provides a detailed visualization of the learning curves for various algorithms in the Jinan scenario on medium traffic. TD3+BC and CQL improve queue length reduction over time, with TD3+BC converging faster than CQL, but they still result in higher overall queue lengths compared to OffLight. OffLight (TD3+BC) and OffLight (CQL) both significantly outperform the other baselines, maintaining the lowest queue lengths throughout training, with OffLight (TD3+BC) performing the best overall. BC, TD3+BC, and CQL, on the other hand, show more variance and slower convergence. This indicates that OffLight is better at focusing on transitions that matter, effectively reducing congestion more consistently. Importance sampling (IS) ensures that transitions generated under different behavior policies are re-weighted based on their relevance to the current policy, correcting for any mismatch between the offline data and the target policy. This adjustment allows OffLight to more effectively utilize the dataset and focus learning on actions that lead to better traffic flow, resulting in faster convergence and significantly lower queue lengths.
5.1.3 Behavior Representation
Figure 7: t-SNE Representation of Actual Data and Learned Trajectories for Hangzhou scenario
Figure 7 visualizes the t-SNE embeddings of the state-action-next state pairs and the learned latent trajectories for the Hangzhou scenario. The t-SNE visualizations reveal clear separation between the controllers in the latent space. This separation is especially evident between expert, random, and rule-based policies, with distinct clusters forming for each type of policy. This demonstrates the effectiveness of Graph-GMVAE in capturing the latent structure of the dataset.
The results also indicate that OffLight is robust to variations in the quality of the dataset. As the proportion of expert data decreases, OffLight continues to outperform baseline methods. This robustness is critical in real-world settings where data quality may vary significantly across different intersections and traffic conditions.
5.2 Ablation Studies
5.2.1 Performance Comparison with Different Levels of Mixing
Figure 8: Performance Comparison of OffLight with Expert and Random Data Mix
In this study, we analyze the impact of mixing varying proportions of data from expert and random policies on the performance of the OffLight framework compared to conventional offline RL algorithms. Specifically, we explore how OffLight performs under different dataset compositions: 33%, 50%, and 67% expert policy data, with the remainder coming from random policies.
Figure 9 presents the performance comparison across different datasets. As the proportion of expert data decreases and random policy data increases, the performance of all algorithms degrades, which is expected. However, OffLight exhibits a more gradual performance decline compared to TD3+BC and CQL. This indicates that OffLight’s IS and RBPS mechanisms are effective in mitigating the negative effects of suboptimal data.
5.2.2 Effectiveness of Improved Sampling Strategies
Figure 9: Ablation Study on CQL (top) and TD3+BC (bottom) with IS and RBPS
We further investigate the contribution of the two components of OffLight: Importance Sampling (IS) and Prioritized Sampling (PS). Both strategies aim to improve learning from offline datasets by addressing different challenges: IS adjusts for the distributional shift between the behavior policy and the target policy, while RBPS focuses the learning on episodes with higher total rewards.
IS consistently outperforms RBPS in all scenarios (Jinan, Hangzhou, Manhattan) and across all demand levels (low, medium, high). The gap between IS and RBPS widens as the traffic demand increases. This is particularly evident in the high traffic demand scenario, where IS achieves up to a 10% improvement in ATT for CQL and 9% improvement for TD3+BC in Manhattan. Medium and high traffic demand scenarios benefit the most from IS, as seen in Hangzhou and Manhattan, where IS outperforms RBPS by several percentage points, especially as congestion increases.
RBPS improves ATT modestly in all scenarios, but its performance is generally lower than IS. In particular, in medium and high traffic demand scenarios, RBPS provides only incremental improvements (around 4-6% in ATT for both CQL and TD3+BC). RBPS performs more consistently across traffic demand levels, but its gains are notably smaller compared to IS, especially in more complex traffic networks like Manhattan.
IS corrects for the distributional shift between the offline dataset (behavior policy) and the target policy, ensuring that transitions more aligned with the target policy are given higher weight. This correction is particularly crucial in high-demand scenarios, where suboptimal data can significantly distort learning. By focusing on more relevant data, IS ensures the model can make more effective decisions, leading to greater improvements in traffic flow, especially in larger, more complex networks like Manhattan. RBPS, while effective in prioritizing episodes with higher rewards, is more sensitive to noise in mixed-quality datasets, especially when suboptimal policies dominate certain episodes. This can lead to smaller performance gains compared to IS, which dynamically adjusts the importance of transitions based on their relevance to the target policy.
5.3 Discussion
In this study, we proposed OffLight, an offline multi-agent reinforcement learning (MARL) framework tailored for traffic signal control (TSC) using heterogeneous, pre-collected datasets. The experiments illustrate that OffLight not only outperforms existing baseline methods but also remains robust across different traffic scenarios and dataset compositions. Key insights from the results include:
Performance with Mixed-Quality Data: OffLight consistently demonstrates superior performance in scenarios involving heterogeneous behavior policies. This is critical for real-world deployments where datasets often comprise both expert and suboptimal policy data. The Graph-GMVAE was crucial in capturing the latent behavior distributions, enabling more accurate policy learning.
Effectiveness of Sampling Strategies: Both importance sampling (IS) and return-based prioritized sampling (RBPS) played pivotal roles in enhancing learning efficiency. IS proved more impactful overall, particularly in high traffic demand scenarios, by correcting for distributional shifts and focusing on transitions more aligned with the target policy. In contrast, RBPS provided more consistent, albeit smaller, gains by prioritizing high-reward episodes. Together, these strategies enabled OffLight to maximize performance, especially in complex environments like Manhattan.
Scalability and Generalization: OffLight scaled well across different network sizes and traffic conditions, from small networks like Jinan to highly complex ones like Manhattan. This demonstrates the framework’s generalizability, making it suitable for deployment in diverse urban settings.
5.3.1 Limitations:
Training Overhead for Graph-GMVAE: Training the Graph-GMVAE, which models the behavior policies, can be computationally expensive, especially in large-scale networks with numerous intersections. This adds to the overall training time and resource requirements, which could limit its real-time deployment capabilities in certain urban environments.
Dependence on Dataset Quality: OffLight’s performance is highly dependent on the quality and diversity of the offline dataset. If the dataset lacks sufficient expert data or is overly biased toward suboptimal policies, OffLight’s learning process may be negatively impacted.
6 Conclusion
OffLight demonstrates the potential to significantly enhance traffic signal control (TSC) in urban networks through offline multi-agent reinforcement learning (MARL). By addressing the key challenges of heterogeneous behavior policies and distributional shifts, OffLight offers a robust solution capable of learning effective policies from mixed-quality datasets. The framework’s incorporation of Graph-GMVAE for behavior policy modeling, alongside importance sampling (IS) and return-based prioritized sampling (RBPS), leads to substantial improvements in traffic efficiency. These advancements reflect the ability of OffLight to scale effectively across different network sizes and traffic conditions, making it a practical tool for offline reinforcement learning for real-world traffic management.
Appendix A OffLight Implementation Details
A.1 Rationale for Averaging Importance Sampling Weights
In multi-agent settings, computing the Importance Sampling (IS) weight as a product across agents can lead to exponential scaling, causing numerical instability and disproportionately amplifying certain transitions. Specifically, the original formulation is given by:
$w_{\text{IS},t}^{i}$ $=\frac{\pi_{\theta}^{i}(A_{t}\mid S_{t})}{\pi_{b}^{i}(A_{t}\mid S_{t})}$ (10)
$=\prod_{i=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid s_{t})}$ (11)
$\approx\prod_{i=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (12)
To mitigate these issues, we adopt an averaging approach:
$w_{\text{IS},t}^{i}$ $\approx\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (13)
Averaging ensures that the combined IS weight remains within a manageable range, preventing the exponential growth associated with multiplication. This approach maintains numerical stability and ensures that each agent’s policy discrepancy contributes proportionally to the overall weight, thereby promoting balanced and stable training dynamics. Empirical results confirm that averaging leads to more stable convergence and prevents any single agent from disproportionately influencing the learning process.
A.2 Pseudocode for Training Behavior Policies
Algorithm 1 summarizes the training procedure of OffLight, highlighting the key steps where Importance Sampling and Return-Based Prioritized Sampling are applied.
Algorithm 1 OffLight Training Procedure
0:  Dataset $\mathcal{D}=\{\tau_{i}\}_{i=1}^{M}$, Target policy parameters $\theta$, Mixer parameters $\phi$
1:  Train GMM-VGAE on $\mathcal{D}$ to model behavior policies
2:  Compute total returns $G_{i}=\sum_{t=0}^{T}r_{t}$ for each trajectory $\tau$
3:  Normalize returns $G_{i}$ to derive RBPS weights $w_{\text{RBPS}}^{i}$
4:  Store estimated behavior policies $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})$ for each episode $i$ in $\mathcal{D}$
5:  Store RBPS weights $w_{\text{RBPS}}^{i}$ for each episode $i$ in $\mathcal{D}$
6:  Initialize target policy parameters $\theta_{\text{target}}\leftarrow\theta$
7:  for each training iteration do
8:     Sample a minibatch of transitions from $\mathcal{D}$ according to RBPS weights $w_{\text{RBPS}}^{i}$
9:     for each transition $(\mathcal{O}_{t}^{i},\mathcal{A}_{t}^{i},\mathcal{R}_{t},\mathcal{O}_{t+1}^{i})$ in the minibatch do
10:        Retrieve stored behavior policy probabilities $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})$ from $\mathcal{D}$
11:        Compute IS weights $w_{\text{IS},t}^{i}$ using:
12:         $w_{\text{IS},t}^{i}=\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$
13:     end for
14:     Combine weights:
15:      $w_{\text{combined},t}^{i}=w_{\text{IS},t}^{i}\times w_{\text{RBPS}}^{i}$
16:     Normalize $w_{\text{combined},t}^{i}$ across the minibatch:
17:      $\tilde{w}_{\text{combined},t}^{i}=\frac{w_{\text{combined},t}^{i}}{\sum_{j}w_{\text{combined},t}^{j}}$
18:     Clamp $\tilde{w}_{\text{combined},t}^{i}$ to prevent extreme values:
19:      $\tilde{w}_{\text{combined},t}^{i}=\text{clamp}(\tilde{w}_{\text{combined},t}^{i},\text{max}=10.0)$
20:     Compute loss $\mathcal{L}$ using:
21:      $\mathcal{L}=\mathbb{E}_{(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\sim\mathcal{D}}\left[\tilde{w}_{\text{combined},t}^{i}\cdot\ell(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\right]$
22:     Update policy parameters $\theta$ using gradient descent on $\mathcal{L}$
23:     Update target policy parameters $\theta_{\text{target}}$ using Polyak averaging or hard updates
24:  end for
A.3 GMM-VGAE Architecture
A.3.1 Encoder: Capturing Spatial and Temporal Dependencies
The encoder processes trajectory data by embedding local observations and actions through linear layers. GATs capture the spatial dependencies among the $N$ traffic intersections by leveraging the graph structure of the traffic network. Subsequently, RNNs, specifically Long Short-Term Memory (LSTM) networks, encode the temporal sequence of these embeddings to capture the action-observation history over time. This dual encoding ensures that both spatial relationships and temporal dynamics are effectively represented in the encoded features.
Formally, the encoding process is defined as: {dmath} μ_t, logσ_t^2 = Encoder(O_t, A_t-1, A), z_t = μ_t + σ_tϵ,  ϵN(0, I), where:
  • $\mathcal{O}$ represents the observations.
  • $\mathcal{A}$ denotes the actions.
  • $\text{Encoder}(\cdot)$ maps the input observations and actions to the parameters of the posterior distribution over the latent variables, yielding the mean $\boldsymbol{\mu}_{t}$ and log-variance $\log\boldsymbol{\sigma_{t}}^{2}$.
  • $\mathbf{z}_{t}$ is the sampled latent variable obtained via the reparameterization trick.
A.3.2 Latent Space: Modeling Behavior Policies
The latent space is structured to represent the distribution of underlying behavior policies using a Gaussian Mixture Model (GMM). By modeling the latent variables $\mathbf{z}$ as a mixture of $K$ Gaussian components, OffLight captures the multimodal nature of behavior policies present in the heterogeneous dataset. This mixture model facilitates the disentanglement of different policy behaviors, enabling accurate estimation and differentiation of effective and suboptimal policies within the data.
The GMM is defined as:
$p(\mathbf{z})$ $=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{z}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}),$ (14)
$q(\mathbf{z}\mid\tau_{i})$ $=\sum_{k=1}^{K}\gamma_{ik}\mathcal{N}(\mathbf{z}\mid\boldsymbol{\mu}_{ik},\boldsymbol{\Sigma}_{ik}),$ (15)
where:
  • $K$ is the number of mixture components.
  • $\pi_{k}$ are the mixture weights with $\sum_{k=1}^{K}\pi_{k}=1$.
  • $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ are the mean and covariance of the $k$-th Gaussian component.
  • $\gamma_{ik}$ represents the responsibility of the $k$-th component for trajectory $\tau_{i}$.
Figure 10: Evaluation of different latent dimensions for modeling behavior policy distribution on Hangzhou scenario
Number of Latent Variables:
The number of latent variables in the Graph-GMM-VGAE model was determined through a validation study using the Hangzhou traffic scenario. We tested several configurations, ranging from 2 to 32 latent dimensions, and measured the reconstruction accuracy of the behavior policies on the validation set. As the number of latent variables increased, the model’s ability to capture heterogeneity in the policy distributions improved. However, beyond 8 latent dimensions, the validation accuracy plateaued, indicating diminishing returns in performance improvement. Additionally, higher-dimensional models risk overfitting, particularly with smaller datasets. Therefore, we selected 8 latent variables as the optimal configuration, offering a good trade-off between accuracy and model complexity.
A.3.3 Decoder: Reconstructing Policy Distributions
The decoder reconstructs the action distributions from the latent variables by concatenating the sampled latent vectors with the encoded observations and passing them through additional GAT layers. This process effectively decodes the policy distributions corresponding to different latent behaviors. The reconstruction ensures that the latent representations accurately capture the decision-making processes of the behavior policies, enabling reliable policy reconstruction and subsequent importance sampling.
The decoding process is formalized as:
$\hat{\mathcal{A}_{t}}$ $=\text{Decoder}(\mathbf{z}_{t},\mathcal{O},\mathcal{A_{t-1}},\mathbf{A}),$ (16)
$p(\mathcal{A}_{t}\mid\mathbf{z}_{t},\mathcal{O}_{t})$ $=\prod_{i=1}^{N}p(a_{t}^{i}\mid\mathbf{z},\mathcal{O}_{t}),$ (17)
where:
  • $\hat{\mathcal{A}}$ represents the reconstructed actions.
  • $\text{Decoder}(\cdot)$ maps the latent variable $\mathbf{z}$ and observations $\mathcal{O}$ back to the action space.
  • $p(a_{t}^{i}\mid\mathbf{z}_{t},\mathcal{O}_{t})$ is modeled as a categorical distribution over possible actions for agent $n$.
OffLight employs a self-supervised learning approach, eliminating the need for explicit labels regarding behavior policies. By minimizing the reconstruction loss alongside the Kullback-Leibler (KL) divergence between the posterior and prior distributions in the latent space, the model learns to encode and decode behavior policies effectively. This self-supervised mechanism ensures that the latent space is structured to accurately reflect the diversity of behavior policies in the dataset.
A.4 Offline RL Framework Details
A.5 Hyperparameters
Component Parameter Value
Encoder LSTM hidden size 128
FC hidden size 128
GAT Layers 3
Attention Heads 4
Latent variable dimensions 8
Hidden layer activation function LeakyReLU
Decoder GAT Layers 3
Attention Heads 4
FC hidden size 128
Hidden layer activation function LeakyReLU
Final layer activation function Softmax
Training Epochs 100
Learning rate 3e-3
Optimizer Adam
Batch size 128
KL divergence weight 1e-4
Table 4: GMM-VGAE Hyperparameters and Architecture Details
Component Parameter Value
Agent Network RNN hidden size 128
FC hidden size 128
GAT Layers 1
Attention Heads 4
Hidden layer activation function ReLU
CQL $\alpha$ 1.0
Target network update interval 200 (episodes)
TD3+BC $\alpha$ 1.0
Target network update $\tau$ 0.005
Training Optimizer Adam
Learning rate 0.001
Table 5: Offline RL Hyperparameters and Architecture Details
We used PyTSC library [28] and the CityFlow simulator [29] to generate traffic scenarios for our experiments, while the OffLight algorithm was implemented using the OffPyMARL framework [30]. Each offline reinforcement learning algorithm was trained for 100 epochs on a system equipped with an Intel(R) Core(TM) i9-9900K CPU @ 3.60GHz and an NVIDIA GeForce RTX 2080 GPU. The entire training process for OffLight took approximately 4 hours per run, with the GMM-VGAE component requiring between 5 to 20 minutes to complete its training.
Appendix B Theoretical Analysis
We analyze the properties of the importance sampling integration. This analysis underscores the robustness and effectiveness of OffLight in offline multi-agent reinforcement learning (MARL) settings, particularly for traffic signal control (TSC) applications.
B.1 Importance Sampling Weight Consistency
In OffLight, importance sampling is integrated to correct for the distributional shift between the behavior policy $\pi_{b}$ and the target policy $\pi_{\theta}$. We aim to show that the importance sampling weights computed using the estimated behavior policy $\hat{\pi}_{b}$ provide consistent and unbiased estimates of expectations under the target policy. For any function $f(\tau)$ of a trajectory $\tau$, the expectation under $\pi_{\theta}$ can be estimated using samples from $\pi_{b}$ as:
$\mathbb{E}_{\tau\sim\pi_{\theta}}[f(\tau)]=\mathbb{E}_{\tau\sim\pi_{b}}\left[w(\tau)f(\tau)\right],$ (18)
where the importance weight $w(\tau)$ is given by:
$w(\tau)=\prod_{t=0}^{T}\prod_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid s_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid s_{t}^{i})}.$ (19)
Assumptions:
  1. 1.
    1. The estimated behavior policy $\hat{\pi}_{b}$ satisfies $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid s_{t}^{i})=\pi_{b}^{i}(a_{t}^{i}\mid s_{t}^{i})$ for all $n,t$.
  2. 2.
    The support of $\pi_{\theta}^{i}(a\mid s)$ is contained within the support of $\pi_{b}^{i}(a\mid s)$ for all $n$, i.e., $\pi_{b}^{i}(a\mid s)=0\implies\pi_{\theta}^{i}(a\mid s)=0$.
Proof: Under these assumptions, the importance weights provide an unbiased estimate:
[…]
$=\mathbb{E}_{\tau\sim\pi_{\theta}}[f(\tau)].$ (23)
This demonstrates that OffLight’s importance sampling weights yield consistent estimates when the behavior policy is accurately estimated.
B.2 Variance Reduction via Accurate Behavior Policy Estimation
While importance sampling provides unbiased estimates, the variance of the estimator can be high, particularly when the target and behavior policies differ significantly. By accurately estimating $\pi_{b}$ using the GMM-VGAE, OffLight reduces the variance of the importance weights.
Variance Analysis: The variance of the importance sampling estimator is given by:
$\text{Var}_{\pi_{b}}[w(\tau)f(\tau)]=\mathbb{E}_{\pi_{b}}[w^{2}(\tau)f^{2}(\tau)]-\left(\mathbb{E}_{\pi_{\theta}}[f(\tau)]\right)^{2}.$ (24)
An accurate estimation of $\pi_{b}$ ensures that $w(\tau)$ does not have extremely large values, which would otherwise inflate the variance. Specifically, when $\pi_{\theta}$ is close to $\pi_{b}$, $w(\tau)$ is close to 1, minimizing the variance.
Theorem: If $\hat{\pi}_{b}$ is a consistent estimator of $\pi_{b}$, then as the amount of data increases, the variance of the importance sampling estimator decreases.
Proof: As $\hat{\pi}_{b}$ converges to $\pi_{b}$, the distribution of $w(\tau)$ becomes tighter around 1, reducing the variance. This is formalized by the Law of Large Numbers and convergence properties of consistent estimators.
B.3 Computational Complexity and Scalability
OffLight leverages Graph Neural Networks (GNNs) to capture spatial dependencies among agents. The computational complexity per forward pass is:
$\mathcal{O}(L\cdot(N+E)\cdot d),$ (25)
where: - $L$ is the number of GNN layers. - $N$ is the number of nodes (agents). - $E$ is the number of edges (inter-agent connections). - $d$ is the dimensionality of the embeddings. In sparse traffic networks, $E=\mathcal{O}(N)$, so the complexity scales linearly with $N$. This demonstrates that OffLight is scalable to large networks, such as the Manhattan scenario with 196 intersections.

### v3

@@
 Abstract
-Efficient traffic signal control (TSC) is essential for modern urban mobility, but traditional systems often struggle to adapt to the complex and dynamic nature of city traffic.
-While Multi-Agent Reinforcement Learning (MARL) offers promising adaptive solutions, online MARL methods require a significant amount of interactions with the environment, which can be expensive and time consuming.
-Offline MARL addresses these concerns by leveraging historical traffic data for training, but it faces challenges due to the heterogeneity of behavior policies in real-world datasets—a mix of different controllers makes learning difficult.
-We introduce OffLight, a novel offline MARL framework specifically designed to handle heterogeneous behavior policies within TSC datasets.
-OffLight employs Gaussian Mixture Model Variational Graph Autoencoder (GMM-VGAEs) to model the complex distribution of behavior policies, enabling effective learning from diverse data sources.
-To enhance coordination between agents, we integrate Graph Attention Networks (GATs), allowing agents to make informed decisions based on aggregated information from neighboring intersections.
-Furthermore, OffLight incorporates Importance Sampling (IS) to correct for differences between the behavior and target policies and utilizes Return-Based Prioritized Sampling (RBPS) to focus on high-quality experiences, thereby improving sample efficiency.
-Extensive experiments across three real-world urban traffic scenarios—Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections)—demonstrate that OffLight significantly outperforms existing offline RL methods.
-Notably, OffLight achieves up to a 7.8% reduction in average travel time and an 11.2% decrease in queue length compared to baseline algorithms, particularly in datasets with mixed-quality data.
-Our ablation studies confirm the effectiveness of OffLight’s components in handling data heterogeneity and enhancing learning performance.
-These results highlight OffLight’s ability to accurately model heterogeneous behavior policies, mitigate the impact of suboptimal data, and scale to large urban networks.
-By addressing the critical challenges of offline MARL in TSC, OffLight offers a practical and impactful solution for improving urban traffic management without the risks associated with online learning.
-Keywords: Multi-Agent Reinforcement Learning, Offline Reinforcement Learning, Traffic Signal Control, Intelligent Transportation Systems, Deep Reinforcement Learning.
-1 Introduction
-Efficient traffic signal control (TSC) is important for modern urban mobility, directly affecting congestion levels, travel times, and overall city livability.
-As urban populations grow and vehicular usage intensifies, traditional traffic management systems struggle to adapt to dynamic and complex traffic flows.
-Recent advancements in Multi-Agent Reinforcement Learning (MARL) offer promising solutions by enabling decentralized, adaptive, and intelligent control of traffic signals [1, 2].
+Efficient traffic signal control (TSC) is crucial for reducing congestion and improving urban mobility.
+While multi-agent reinforcement learning (MARL) has shown promise in adaptive traffic management, its reliance on real-time interactions makes online training costly and impractical.
+Offline MARL, which learns from historical data, offers a safer and more scalable alternative but struggles with heterogeneous behavior policies in real-world datasets.
+We introduce OffLight, a novel offline MARL framework designed to handle policy heterogeneity in traffic signal control datasets.
+OffLight integrates importance sampling (IS) to correct for distributional shifts and return-based prioritized sampling (RBPS) to emphasize high-quality experiences.
+To model diverse behavior policies, it leverages a Gaussian mixture model variational graph autoencoder (GMM-VGAE), capturing spatial and temporal traffic dynamics while improving policy estimation.
+We evaluate OffLight across real-world urban traffic scenarios varying from small to large traffic signal networks.
+Results show that OffLight outperforms existing offline RL methods, reducing average travel time by up to 7.8% and queue length by 11.2% compared to state-of-the-art approaches.
+Unlike prior offline RL methods, OffLight effectively adapts to complex and mixed-policy datasets, making it more reliable for real-world deployment without risky online training.
+These results highlight OffLight’s potential to improve urban traffic management at scale.
+Our implementation is available at https://github.com/rbokade/offlight.
+I INTRODUCTION
+Efficient traffic signal control (TSC) is critical for urban mobility, directly affecting congestion, travel times, and city livability.
+Multi-agent reinforcement learning (MARL) has emerged as a promising solution by enabling decentralized, adaptive, and intelligent control of traffic signals [1, 2].
+While online MARL methods rely on real-time interaction, offline MARL uses pre-collected data, offering safer, more cost-efficient, and scalable alternatives by eliminating the need for risky real-time experimentation [3, 4, 5, 6].
 Figure 1: General Offline MARL Framework for Traffic Signal Control
-While online MARL requires agents to interact with the environment in real-time—which can be impractical and unsafe in real-world traffic systems—offline MARL leverages historically collected traffic data to train agents without live experimentation [3].
-This approach offers several advantages:
-  • Extensive Data Utilization: Utilizes pre-collected traffic data, allowing agents to learn from diverse scenarios and rare events difficult to replicate online [4].
-  • Safety and Risk Mitigation: Eliminates risks associated with deploying untested policies in live traffic, ensuring only well-trained agents are implemented [5].
-  • Cost Efficiency and Scalability: Reduces the need for expensive simulations or field tests, facilitating deployment across various scales and environments [3].
-  • Accelerated Policy Development: Enables rapid iteration and refinement of control policies, expediting research and development [6].
-Several offline reinforcement learning (RL) methods have been explored for TSC, often incorporating imitation learning to bootstrap training.
-For instance, DemoLight uses demonstrations from traditional controllers to initialize policies, enabling efficient exploration and faster convergence compared to pure RL approaches [7].
-Cooperative Control for multi-intersection TSC combines imitation learning and deep RL to coordinate multiple intersections, highlighting the benefits of learning from expert demonstrations [8].
-CrossLight begins with offline training using behavior cloning and transitions to online RL to adapt policies to new environments, showcasing the effectiveness of imitation learning for strong policy initialization [9].
-Building upon these efforts, DataLight introduces an input space modeling technique that captures the spatial distribution of traffic using self-attention mechanisms, demonstrating improved policy performance without requiring live interactions [4].
-Figure 1 represents a general framework of an offline RL algorithm.
-However, a key challenge in offline MARL for TSC remains unaddressed: real-world datasets often consist of a heterogeneous mix of behavior policies, making it difficult for agents to learn effective strategies.
-This heterogeneity can arise from a variety of factors, such as:
-  • Local Heuristics: Many traffic signals operate on traditional rule-based controllers (e.g., fixed-time or reactive control).
-These heuristics are often suboptimal and vary between intersections [10].
-  • Varying Control Strategies: Over time, traffic systems may switch between different controllers (e.g., from manual control to AI-based systems), resulting in datasets that reflect a blend of expert and suboptimal policies [11].
-  • Temporal Variability: Traffic patterns fluctuate throughout the day, leading to variations in the policies being implemented during different time periods.
-For instance, a traffic signal might switch between morning peak-time strategies and off-peak strategies [10].
-This heterogeneity hampers effective learning because:
-  • Offline MARL methods are highly sensitive to the underlying distribution of policies in the dataset.
-Mixing high- and low-quality data complicates the learning process, leading to poor generalization [12].
-  • Mixed-policy data can introduce biases, making it harder for the agent to identify the most effective policies [13].
-To address these challenges, we introduce OffLight, a novel offline MARL framework that combines Importance Sampling (IS) and Return-Based Prioritized Sampling (RBPS) to mitigate distributional shifts and focus on high-value experiences.
-This synergy between IS and RBPS enhances both sample efficiency and policy performance.
-GMM-VGAE models the diversity in behavior policies within the traffic network, capturing policy heterogeneity by disentangling them in the latent space.
-By leveraging Graph Neural Networks (GNNs), OffLight combines local observations across intersections into a structured global representation, which is crucial in traffic signal control where agent interactions are highly localized but collectively influence network-wide patterns.
-This enables OffLight to use IS effectively, correcting for distributional shifts and ensuring stable policy learning from varied data.
-Meanwhile, RBPS prioritizes high-return episodes, allowing OffLight to focus on the most informative experiences, thus accelerating convergence and optimizing performance.
-Key contributions of this paper include:
-  • Developing OffLight, an offline MARL framework designed to handle heterogeneous behavior policies in traffic signal control (TSC) using a combination of Importance Sampling (IS) and Return-Based Prioritized Sampling (RBPS).
-  • Introducing GMM-VGAE to accurately capture diverse behavior policies, supporting robust learning under policy heterogeneity and enabling effective IS.
-  • Leveraging RBPS to prioritize high-return episodes, which enhances sample efficiency and accelerates convergence by focusing on the most informative experiences.
-  • Demonstrating OffLight’s scalability and superior performance on real-world traffic datasets, where it consistently outperforms existing methods.
-The remainder of the paper is organized as follows.
-Section 2 reviews related work and provides necessary background on MARL and offline RL in the context of TSC.
-Section 3 details the OffLight framework, including its architecture and key components.
-Section 4 describes the experimental setup, and Section 5 presents the results and analysis.
-Finally, Section 6 concludes the paper and Section LABEL:sec:discussion discusses the key insights gained from this study and limitations of OffLight.
-2 Background
-2.1 Multi-Agent Reinforcement Learning
-Multi-Agent Reinforcement Learning (MARL) involves multiple agents interacting within a shared environment.
-Each agent learns to maximize the cumulative reward while adapting to the behaviors of other agents.
-MARL introduces additional challenges such as non-stationarity (due to concurrent learning), the need for coordination, and partial observability, all of which make policy learning significantly harder than in single-agent settings [14, 15].
-2.1.1 Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs)
-Multi-agent problems are commonly modeled as Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs) [16].
-A Dec-POMDP is formally defined by a tuple $\langle N,\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{N},\{\mathcal{O}^{i}\}_{i=1}^{N},P,R,H,\gamma\rangle$, where:
-  • $N$ is the number of agents.
-  • $\mathcal{S}$ is the set of environment states.
-  • $\mathcal{A}^{i}$ is the action space of agent $i$.
-  • $\mathcal{O}^{i}$ is the observation space of agent $i$.
-  • $P:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is the state transition probability function.
-  • $R:\mathcal{S}\times\mathcal{A}^{1}\times\dots\times\mathcal{A}^{N}\rightarrow\mathbb{R}$ is the shared reward function.
-  • $H$ is the finite horizon.
-  • $\gamma\in[0,1)$ is the discount factor.
-Each agent $i$ selects actions based on its action-observation history $h^{i}_{t}=\{o^{i}_{0},a^{i}_{0},\dots,o^{i}_{t-1},a^{i}_{t-1},o^{i}_{t}\}$, where $o^{i}_{t}\in\mathcal{O}^{i}$ is the observation at time $t$, and $a^{i}_{t}\in\mathcal{A}^{i}$ is the action.
-The joint action $\mathbf{a}_{t}=(a_{t}^{1},\dots,a_{t}^{N})$ determines the next state $s_{t+1}$ according to $P$, and all agents receive a shared reward $R$.
-The objective in Dec-POMDPs is to find a set of decentralized policies $\{\pi^{i}\}_{i=1}^{N}$, where each $\pi^{i}:H^{i}_{t}\rightarrow\mathcal{A}^{i}$ maps action-observation histories to actions, such that the expected cumulative reward is maximized:
+Several offline RL methods have shown promise in traffic signal control by using imitation learning to bootstrap initial training.
+Methods such as DemoLight [7], Cooperative Control [8], and CrossLight [9] have leveraged expert demonstrations to quickly initialize and refine traffic control policies.
+More recently, DataLight [4] improved policy performance through attention mechanisms that capture spatial traffic distributions.
+However, these approaches do not fully address a significant issue in offline MARL datasets: heterogeneous behavior policies arising from varied control strategies, temporal variability, and the mixture of different operational contexts [10, 11].
+The heterogeneity in real-world traffic datasets poses significant challenges for offline MARL, primarily due to distributional shifts and suboptimal behavior biases that degrade policy learning [12, 13].
+Recent studies, including Hong et al.[14] and Yue et al.[15], have proposed prioritized resampling techniques to emphasize high-quality experiences, thus improving policy robustness in offline settings.
+Nevertheless, effectively modeling diverse behavior policies in complex multi-agent environments remains a substantial challenge for offline MARL methods.
+To address these challenges, we propose OffLight, an offline MARL framework specifically designed for handling heterogeneous behavior policies in TSC datasets.
+OffLight integrates importance sampling (IS) to correct distributional shifts and return-based prioritized sampling (RBPS) to highlight high-quality experiences.
+Additionally, OffLight uses a Gaussian mixture model variational graph autoencoder (GMM-VGAE) to represent diverse behavior policies accurately.
+Experiments across real-world urban traffic scenarios show that OffLight outperforms existing offline RL methods, demonstrating reduced average travel times and queue lengths.
+Our results highlight OffLight’s capability to reliably manage urban traffic networks without the risks associated with online training, making it a practical and scalable solution for real-world deployment.
+II Background
+II-A Multi-Agent Reinforcement Learning
+We model the MARL problem as Decentralized Partially Observable Markov Decision Processes (Dec-POMDPs) [16], formally defined by the tuple $\langle N,\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{N},\{\mathcal{O}^{i}\}_{i=1}^{N},P,R,H,\gamma\rangle$.
+Here, $N$ is the number of agents, $\mathcal{S}$ is the state space, $\mathcal{A}^{i}$ and $\mathcal{O}^{i}$ are the action and observation spaces for agent $i$, respectively, $P$ is the state transition probability, $R$ is the shared reward, $H$ is the finite horizon, and $\gamma\in[0,1)$ is the discount factor.
+Each agent $i$ selects actions based on its local action-observation history $h^{i}_{t}=\{o^{i}_{0},a^{i}_{0},\dots,o^{i}_{t}\}$, resulting in joint action $\mathbf{a}_{t}=(a_{t}^{1},\dots,a_{t}^{N})$, transitioning the environment to state $s_{t+1}$ according to $P$, and yielding a shared reward $R$.
+The objective is to determine decentralized policies $\{\pi^{i}\}_{i=1}^{N}$, with each policy $\pi^{i}:H^{i}_{t}\rightarrow\mathcal{A}^{i}$ mapping histories to actions, to maximize expected cumulative reward [^1] In practice, we employ local rewards for each agent and adopt Centralized Training with Decentralized Execution (CTDE), as we found this approach converges faster than using a scalar global reward.:
 $\max_{\{\pi^{i}\}}\mathbb{E}\left[\sum_{t=0}^{H}\gamma^{t}R(s_{t},\mathbf{a}_{t})\right].$ (1)
-2.2 Application to Traffic Signal Control (TSC)
-In the context of Traffic Signal Control (TSC), each traffic signal controller can be modeled as an agent within the Decentralized Partially Observable Markov Decision Process (Dec-POMDP) framework [16].
-Each agent operates independently and makes decisions based on local observations, such as vehicle counts, vehicle speeds, queue lengths, and current signal phases.
-Observation Space: Each traffic signal agent has a limited range of observation.
-This range reflects the realistic coverage of common sensors such as inductive loop detectors or cameras.
-For each incoming lane, the agent observes:
-  • The number of vehicles $\{n_{l}\}_{l=1}^{L_{i}}$, where $L_{i}$ is the number of lanes for traffic signal $i$.
-  • The average speed of vehicles $\{s_{l}\}_{l=1}^{L_{i}}$, normalized by the speed limit.
-  • The number of halted vehicles, representing the queue length $\{q_{l}\}_{l=1}^{L_{i}}$.
-  • The current phase ID, indicating the traffic signal’s current phase.
-These local observations provide the sensory input required for each agent to make real-time decisions about the traffic signal phases.
-Action Space: Each traffic signal agent controls the signal phases at its intersection.
-The available actions correspond to selecting one of the predefined green phases from the list of possible phases for that intersection.
-An agent may select a green phase and it must then transition to the yellow phase, which is enforced by the environment.
-The action selection interval is fixed at 5 simulation seconds.
-Reward Function: A commonly used reward metric is queue length.
-Minimizing the queue length across all lanes in the network directly reflects the reduction of vehicle stoppage.
-The global reward at each time step $r_{t}$ is defined as:
-$r_{t}=-\sum_{l\in L}q_{l}(t),$
-where $q_{l}(t)$ represents the queue length on lane $l$ at time $t$, and $L$ is the set of all lanes in the network.
-This reward structure encourages agents to minimize congestion and improve traffic flow throughout the system.
-Deep MARL: In multi-agent TSC, the actions of one traffic signal directly impact the observations and performance of neighboring intersections.
-Therefore, coordination between agents is crucial to achieve optimal traffic management [17].
-Traditional approaches, such as Independent Q-Learning (IQL) [18], treat each agent as independent, which may lead to suboptimal coordination.
-Recent advancements in the application of deep MARL in TSC include frameworks such as CoLight [11], PressLight [19], AttendLight [20] which utilize Graph Attention Networks (GATs) [21, 11] to model inter-agent communication.
-These frameworks model the traffic network as a graph, where intersections are treated as nodes and roads as edges, allowing traffic signals to share information with neighboring intersections.
-By focusing attention on the most relevant neighbors, these methods have achieved significant success in optimizing traffic flow through coordinated agent behavior.
-Following the success of these frameworks in the online MARL setting, we choose this approach as the foundation for our framework to model offline MARL policies.
-By leveraging the well-established GAT-based communication mechanism, we aim to maintain effective coordination between agents, even when learning occurs from static offline datasets.
-This approach ensures that OffLight can efficiently handle the complexity of real-world traffic systems, while addressing the challenges inherent in offline learning.
-2.3 Offline Multi-Agent Reinforcement Learning
-Offline Multi-Agent Reinforcement Learning involves learning effective policies from a fixed dataset of interactions without further environmental exploration during training [3, 4].
-This setting is crucial in scenarios where data collection is expensive, risky, or impractical, such as traffic signal control.
-Offline MARL introduces unique challenges:
-  • Distributional Shift: The static dataset may not adequately represent the state-action distribution encountered when executing the learned policies, leading to overestimation of out-of-distribution actions and degraded performance [22].
-  • Extrapolation Error: Without any interaction with the environment, agents must generalize based solely on the dataset provided.
-This limitation risks extrapolation errors, where agents might attempt actions unsupported by the data, leading to poor or unsafe decisions when deployed.
-  • Limited Coverage of Optimal Strategies: Offline datasets may lack the comprehensive coverage needed for optimal joint policies.
-This is particularly challenging in MARL, as optimal coordination patterns often require data from multiple interacting policies, which may be missing in historical data.
-Recent offline RL algorithms, such as Conservative Q-Learning (CQL) [12] and TD3+BC [23], have been developed to address these challenges by regularizing learned policies and penalizing overestimation of Q-values for unseen actions.
-While originally designed for single-agent settings, these methods can be extended to multi-agent systems using frameworks like Centralized Training with Decentralized Execution (CTDE) [24].
-2.3.1 Conservative Q-Learning (CQL)
-Conservative Q-Learning (CQL) [12] is an offline RL algorithm designed to address the overestimation of out-of-distribution (OOD) actions by learning conservative Q-value estimates.
-CQL modifies the standard Bellman error minimization by adding a regularization term that penalizes Q-values for actions not present in the dataset.
-The CQL objective is: {dmath} min_Q α( E_s ∼D, a ∼π(a|s) [Q(s,a)] - E_(s,a) ∼D [Q(s,a)] ) + L_Bellman(Q), where $\mathcal{D}$ is the offline dataset, $\pi(a|s)$ is a policy (e.g., the uniform policy), $\alpha$ is a regularization coefficient, and $\mathcal{L}_{\text{Bellman}}(Q)$ is the standard Bellman error.
-In multi-agent settings, the joint action space grows exponentially with the number of agents, making it infeasible to represent the full joint Q-function.
-To address this, CQL can be extended using factorization methods or the CTDE paradigm [24].
-CTDE allows agents to be trained centrally with access to global information, while at execution time, they act based on local observations, enabling scalability and coordination among agents even with limited data.
-2.3.2 TD3+BC
-TD3+BC [23] enhances the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm [25] by integrating a behavior cloning (BC) term into the policy update.
-This addition encourages the learned policy to stay close to the behavior policy present in the offline dataset, mitigating issues arising from distributional shift.
-The modified policy objective in TD3+BC is:
-$\min_{\pi_{\theta}}-\mathbb{E}_{s\sim\mathcal{D}}[Q_{\phi}(s,\pi_{\theta}(s))]+\lambda\mathbb{E}_{(s,a)\sim\mathcal{D}}[\|\pi_{\theta}(s)-a\|_{2}^{2}],$ (2)
-where $\pi_{\theta}$ is the policy network parameterized by $\theta$, $Q_{\phi}$ is the critic network parameterized by $\phi$, and $\lambda$ balances the importance of the behavior cloning term.
-In multi-agent systems, TD3+BC can be adapted by allowing each agent to learn its own policy while considering the interactions with other agents.
-Using the CTDE framework [26], agents are trained with access to additional global information (e.g., other agents’ actions or policies) to improve coordination.
-During execution, agents operate based on their local observations, ensuring scalability and feasibility in decentralized environments.
-2.4 Challenge of Heterogeneous Behavior Policies in Offline MARL
-A unique challenge in offline MARL, particularly in domains like Traffic Signal Control (TSC), is managing heterogeneous behavior policies present in the offline dataset.
-In TSC, data often come from various sources, times, or operational contexts, leading to a mix of behavior policies that agents must learn from.
-This heterogeneity adds complexity to offline MARL in the following ways:
-  • Increased Risk of Distributional Shift: When policies in the offline data vary significantly, the likelihood of a distributional mismatch between the training data and the learned policy increases.
-The agents are more prone to encountering out-of-distribution actions, resulting in unreliable Q-value estimates and degraded policy performance, as the diversity in behavior data can make distributional correction methods less effective [22].
-  • Difficulty in Behavior Policy Estimation: For methods relying on importance sampling or policy constraints, accurately estimating a single behavior policy becomes challenging when the data reflects multiple, potentially conflicting policies.
-Modeling this mixture accurately is essential to effective learning but is significantly more complex in offline MARL with heterogeneous datasets [13].
-Addressing the challenge of heterogeneous behavior policies is essential for effective offline MARL, especially in domains like TSC where operational conditions and policies can vary widely.
-2.5 Motivation for the OffLight Framework
-To tackle the specific challenge of heterogeneous behavior policies in offline MARL, we propose the OffLight framework. OffLight is designed to:
-  • Model Heterogeneous Behavior Policies: Utilize advanced modeling techniques, such as Variational Graph Auto-Encoders (GMM-VGAEs), to accurately estimate and represent the diverse behavior policies present in the offline dataset.
-  • Mitigate Distributional Shift: Employ importance sampling with the estimated behavior policy to correct for discrepancies between the behavior policies and the target policy, reducing overestimation and improving policy evaluation.
-  • Improve Sample Efficiency: In addition to IS and GMM-VGAE, OffLight integrates Return-Based Prioritized Sampling (RBPS) based on episodic returns.
-RBPS prioritizes high-reward episodes, ensuring that the learning process focuses on the most informative and successful experiences.
-This prioritization accelerates convergence and enhances the robustness of the learned policies.
-By focusing on accurately modeling and leveraging the heterogeneous behavior policies in the offline data, OffLight addresses a critical challenge in offline MARL for TSC.
-3 OffLight Framework
-OffLight is designed to tackle the challenges of offline multi-agent reinforcement learning (MARL) in the domain of traffic signal control (TSC).
-It combines a self-supervised learning approach with importance sampling (IS) to account for heterogeneous policies and improve learning from static, pre-collected datasets.
-IS and RBPS weights are collected before the offline RL training and stored.
+To optimize decentralized policies, Deep Q-Learning (DQNs) [17] is commonly employed, where each agent independently updates its Q-values based on local observations and rewards as follows: {dmath} Q^i(o^i_t,a^i_t) ←Q^i(o^i_t,a^i_t) + α(r^i_t + γmax_a’ Q^i(o^i_t+1,a’) - Q^i(o^i_t,a^i_t)).
+However, independent learning may lead to suboptimal coordination.
+Recent frameworks, such as CoLight [11], PressLight [18], and AttendLight [19], address this by incorporating Graph Attention Networks (GATs) [20] for inter-agent communication.
+We leverage these successful frameworks as the basis for our approach.
+II-B Offline Multi-Agent Reinforcement Learning
+Offline Multi-Agent Reinforcement Learning (Offline MARL) learns policies from fixed datasets without additional environment interaction during training [3, 4].
+Offline MARL presents several challenges: (1) Distributional Shift: The dataset may differ significantly from the policy’s actual environment interactions, leading to inaccurate value estimations and reduced performance [21].
+(2) Extrapolation Error: Without interactive feedback, agents risk generalizing poorly to situations not covered by the dataset.
+(3) Limited Coverage: Historical data might not include optimal coordination strategies necessary for effective multi-agent policies.
+Recent offline RL algorithms address these issues.
+Conservative Q-Learning (CQL) [12] introduces a penalty for overestimating Q-values of actions not in the dataset.
+Its objective is: {dmath} min_Q_i α( Eo_i ∼D, a_i ∼π(a_i|o_i) [Q(o_i,a_i)] - E(o_i,a_i) ∼D [Q(o_i, a_i)] ) + L_Bellman(Q_i).
+TD3+BC [22] combines TD3 with a behavior cloning term to reduce policy deviation from the dataset, thus managing distributional shift: {dmath} min_π_θ_i -Eo_i ∼D [Qϕ(o_i, π_θ_i(o_i))] + λE(o_i,a_i) ∼D [|π_θ_i(o_i) - a_i|_2^2].
+Both methods can be adapted to multi-agent systems using Centralized Training with Decentralized Execution (CTDE) [23, 24], enabling centralized training and decentralized agent execution for better scalability and coordination.
+II-C Motivation for the OffLight Framework
+Offline MARL faces unique challenges when datasets contain heterogeneous behavior policies, especially in domains like Traffic Signal Control (TSC).
+Different operational contexts or sources create diverse behaviors, increasing the likelihood of distributional mismatch.
+This diversity leads to unreliable Q-value estimates and degraded performance, as distributional correction methods become less effective [21].
+Additionally, accurately estimating mixed behavior policies is difficult, complicating policy constraints and importance sampling [13].
+To address heterogeneous behavior policies in offline MARL, we propose the OffLight framework.
+OffLight utilizes a behavior policy estimation technique (GMM-VGAE) to accurately model policy diversity, employs importance sampling to mitigate distributional shift, and integrates Reward-Based Prioritized Sampling (RBPS) for improved sample efficiency.
+RBPS prioritizes high-reward episodes, enhancing the focus on valuable training data.
+These strategies aim to improve policy robustness and performance specifically in Traffic Signal Control (TSC) scenarios.
+The OffLight framework details are introduced in the next section.
+III OffLight Framework
+OffLight is designed to tackle the challenges of offline MARL in the domain of traffic signal control (TSC).
+It combines a self-supervised learning approach with IS to account for heterogeneous policies and improve learning from static, pre-collected datasets.
+IS and RBPS weights are collected before the offline RL training and stored (see Figure 2).
 During the offline RL training phase, these weights are used to improve learning.
@@
 Figure 2: OffLight Architecture: Integrating GMM-VGAE with IS and RBPS for offline MARL in traffic signal control
-3.1 Dataset Structure
+III-A OffLight Agent Design
+Observation Space: Each traffic signal agent has a limited range of observation.
+This range reflects the realistic coverage of common sensors such as inductive loop detectors or cameras.
+For each incoming lane, the agent observes: (1) number of vehicles $\{n_{l}\}_{l=1}^{L_{i}}$, where $L_{i}$ is the number of lanes for traffic signal $i$, average speed of vehicles $\{s_{l}\}_{l=1}^{L_{i}}$, normalized by the speed limit, number of halted vehicles, representing the queue length $\{q_{l}\}_{l=1}^{L_{i}}$, and the current phase ID, indicating the traffic signal’s current phase.
+These local observations provide the sensory input required for each agent to make real-time decisions about the traffic signal phases.
+Action Space: Each traffic signal agent is responsible for managing the signal phases at its respective intersection.
+The agent’s available actions involve selecting one of the predefined green phases applicable to that intersection.
+Once a green phase is chosen, the system enforces a mandatory transition to a yellow phase.
+The action selection occurs at fixed intervals of 5 simulation seconds.
+Reward Function: We use minimizing the queue length as objective.
+The reward for each agent $i$ at each time step $r^{i}_{t}$ is defined as $r^{i}_{t}=-\sum_{l\in L_{i}}q_{l}(t),$ where $q_{l}(t)$ represents the queue length on lane $l$ at time $t$, and $L_{i}$ is the set of incoming lanes for traffic signal $i$.
+This reward structure encourages agents to minimize congestion and improve traffic flow throughout the system.
+III-B Dataset Structure
 We consider a dataset $\mathcal{D}$ collected from a network of $N$ intersections over $M$ episodic trajectories.
@@
 $\mathcal{D}=\left\{\tau_{k}\right\}_{k=1}^{M},\quad\text{where}\quad\tau_{i}=\left\{\left(\mathcal{O}_{t},\mathcal{A}_{t},\mathcal{O}^{\prime}_{t},\mathcal{R}_{t}\right)\right\}_{t=0}^{T_{k}},$ (2)
-where:
-  • $\mathcal{O}_{t}=\left\{o_{t}^{1},o_{t}^{2},\dots,o_{t}^{N}\right\}$ denotes the observations at time step $t$ for each of the $N$ intersections.
-  • $\mathcal{A}_{t}=\left\{a_{t}^{1},a_{t}^{2},\dots,a_{t}^{N}\right\}$ represents the joint actions taken by the agents at time $t$.
-  • $\mathcal{O}^{\prime}_{t}=\left\{o_{t+1}^{1},o_{t+1}^{2},\dots,o_{t+1}^{N}\right\}$ are the subsequent observations after taking actions $\mathcal{A}_{t}$.
-  • $\mathcal{R}_{t}$ is the reward received at time $t$.
-  • $T_{k}$ is the length of trajectory $\tau_{k}$.
+where $\mathcal{O}_{t}=\left\{o_{t}^{1},o_{t}^{2},\dots,o_{t}^{N}\right\}$ denotes the observations at time step $t$ for each of the $N$ intersections, $\mathcal{A}_{t}=\left\{a_{t}^{1},a_{t}^{2},\dots,a_{t}^{N}\right\}$ represents the joint actions taken by the agents at time $t$, $\mathcal{O}^{\prime}_{t}=\left\{o_{t+1}^{1},o_{t+1}^{2},\dots,o_{t+1}^{N}\right\}$ are the subsequent observations after taking actions $\mathcal{A}_{t}$, $\mathcal{R}_{t}$ is the reward received at time $t$, and $T_{k}$ is the length of trajectory $\tau_{k}$.
 This structure reflects real-world data collected from sensor-equipped intersections and enables learning traffic control policies from historical data.
-3.2 Self-Supervised Learning of Behavior Policies
-To effectively model diverse behavior policies in offline traffic signal control (TSC) data, OffLight uses a Gaussian Mixture Model-based Variational Graph Autoencoder (GMM-VGAE), designed to capture both spatial and temporal dependencies while addressing policy heterogeneity across intersections and time.
-  • GMM for Policy Heterogeneity: The GMM structure in the VGAE models the latent space as a mixture of Gaussian distributions, allowing OffLight to represent multiple underlying policies.
-This disentanglement of varied traffic control behaviors within the latent space enables OffLight to more accurately capture and leverage diverse policies from real-world datasets.
-  • Graph Attention Networks (GATs) and LSTM for Spatial-Temporal Dynamics: GATs capture spatial dependencies by allowing each agent to attend to relevant neighboring intersections.
-This attention mechanism is structured to combine local information effectively, creating a global latent representation of behavior policies across the network.
-Additionally, GATs enable efficient parameter sharing across agents, improving scalability for TSC in a manner similar to MARL by processing local data in a structured and computationally manageable way.
-LSTM layers capture the temporal evolution of traffic patterns by processing sequences of observations over time, preserving the dynamic dependencies needed for effective policy modeling.
-Together, these components enable GMM-VGAE to accurately model both local and global interactions in the traffic network, ensuring OffLight’s robustness to the diverse and dynamic nature of traffic conditions.
-This architecture makes OffLight highly scalable and effective in learning from heterogeneous TSC datasets.
-By constructing a latent space with Gaussian mixtures and capturing spatial-temporal dependencies through GAT and LSTM layers, OffLight provides a structured way to estimate behavior policies at each intersection.
-This enables seamless integration of importance sampling (IS) and Return-Based Prioritized Sampling (RBPS).
-Together, these techniques enhance learning from offline datasets by prioritizing transitions aligned with the target policy and derived from high-reward episodes.
-Below, we discuss how IS and RBPS are applied within OffLight to balance distributional shifts and improve sample efficiency, advancing offline reinforcement learning for traffic signal control (TSC).
-Encoder: Captures spatial dependencies using Graph Attention Networks (GATs) and temporal dynamics with a Long Short Term Memory (LSTM) layer.
-This allows the model to effectively process both the local and global traffic behavior.
-$\mathbf{z}=\text{Encoder}(\mathcal{O}_{t},\mathcal{A}_{t-1},\mathbf{A})$
-Latent Space: Represents the underlying policies as a mixture of Gaussian variables, enabling the disentanglement of different policy behaviors across the network.
-$p(\mathbf{z})=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{z}\mid\mu_{k},\Sigma_{k})$
-Decoder: Reconstructs the policy distribution based on the latent representation and current observation, providing an estimate of the behavior policy for each traffic signal.
-$\mathcal{A}_{t}=\text{Decoder}(\mathcal{O}_{t},\mathbf{z}_{t},\mathbf{A})$
-represents the estimated probability distribution $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\cdot)$ over actions for agent $n$ at time step $t$, conditioned on the current observation and global latent policy representation.
-A benefit of using GATs is that parameters can be shared across agents, thereby reducing model complexity and improving the learning efficiency.
-By learning these latent representations, OffLight can accurately model heterogeneous policies and ensure that the learning process accounts for diverse traffic control strategies.
-A detailed explanation of GMM-VGAE can be found in Appendix A.3.
-3.3 Importance sampling Integration
-To address the challenge of distributional shift between the behavior policy and the target policy, OffLight integrates importance sampling (IS).
+III-C Self-Supervised Learning of Behavior Policies
+OffLight employs GMM-VGAE to model heterogeneous policies, using: (1) GMM-VGAE for Policy Diversity: A Gaussian mixture graph variational autoencoder model (GMM) captures distinct traffic control policies.
+(2) GATs and LSTMs for Spatial-Temporal Dynamics: Graph attention networks (GATs) encode local interactions, while LSTMs track temporal trends.
+These components construct a structured latent space for estimating policies at each intersection.
+The encoder processes agent observations and past actions to capture spatial-temporal dependencies.
+The latent space represents diverse traffic control policies using a GMM-VGAE model.
+Finally, the decoder reconstructs policy distributions based on the latent features and current observations.
+We also enable parameter sharing among agent, reducing model complexity while improving efficiency.
+This design allows OffLight to effectively model heterogeneous policies across intersections.
+III-D Importance sampling Integration
+To address the challenge of distributional shift between the behavior policy and the target policy, OffLight integrates IS.
 This mechanism adjusts the influence of each transition based on its alignment with the target policy, ensuring that the learning algorithm emphasizes relevant and high-quality data.
 $w_{\text{IS},t}^{k}=\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (3)
-where:
-  • $w_{\text{IS},t}^{k}$ is the importance sampling weight for the transition at time step $t$ in episode $k$.
 […]
-  • $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})$ represents the estimated behavior policy probability for agent $n$, obtained from the GMM-VGAE, conditioned on it’s hidden state $\mathbf{h}_{t}^{i}$ and the latent embedding of behavior policy $\mathbf{z}_{t}$.
-Importance sampling corrects for the discrepancy between the behavior policy and the target policy by re-weighting transitions  [^2] The target policy is the policy that the offline RL algorithm seeks to optimize, while the behavior policy refers to the policy used to generate the offline dataset..
+IS corrects for the discrepancy between the behavior policy and the target policy by re-weighting transitions  [^3] The target policy is the policy that the offline RL algorithm seeks to optimize, while the behavior policy refers to the policy used to generate the offline dataset..
 Transitions where the actions are more probable under the target policy relative to the behavior policy receive higher importance sampling weights.
-3.4 Return-Based Prioritized Sampling
-To enhance sample efficiency and accelerate the learning process, OffLight employs Return-Based Prioritized Sampling (RBPS), similar to [27].
+III-E Return-Based Prioritized Sampling
+To enhance sample efficiency and accelerate the learning process, OffLight employs RBPS, similar to [15].
 This strategy prioritizes episodes based on their cumulative rewards, ensuring that the learning algorithm focuses on more successful traffic control experiences.
 $w_{\text{RBPS}}^{k}=C\left(\frac{G_{k}-G_{\text{min}}}{G_{\text{max}}-G_{\text{min}}}+p_{\text{base}}\right)$ (4)
-where:
-  • $w_{\text{RBPS}}^{k}$ is the Return-Based Prioritized Sampling weight for episode $k$.
-  • $G_{k}=\sum_{t=0}^{T_{k}}r_{t}$ is the total reward for episode $k$.
-  • $G_{\text{min}}$ and $G_{\text{max}}$ are the minimum and maximum total rewards across all episodes in the dataset, respectively.
-  • $p_{\text{base}}$ is a small positive constant added to ensure that all episodes have a non-zero probability of being sampled.
-  • $C$ is a normalization constant to maintain numerical stability or ensure the weights sum to a desired value.
+where $w_{\text{RBPS}}^{k}$ is the Return-Based Prioritized Sampling weight for episode $k$, $G_{k}=\sum_{t=0}^{T_{k}}r_{t}$ is the total reward for episode $k$, $G_{\text{min}}$ and $G_{\text{max}}$ are the minimum and maximum total rewards across all episodes in the dataset, respectively.
+$p_{\text{base}}$ is a small positive constant added to ensure that all episodes have a non-zero probability of being sampled, and $C$ is a normalization constant to maintain numerical stability or ensure the weights sum to a desired value.
 RBPS ensures that episodes with higher returns, indicative of more effective traffic management, are sampled more frequently.
@@
 This especially proves beneficial when the distribution of the data is skewed or multimodal as can be seen in Figure 3.
-(a) Jinan
-(b) Hangzhou
-(c) Manhattan
 Figure 3: Distribution of Episodic Returns in mixed-policy datasets.
@@
 RBPS targets this imbalance by prioritizing episodes with higher returns, ensuring that learning is focused on successful traffic control strategies.
-3.5 Combining importance sampling and Return-Based Prioritized Sampling
+III-F Combining importance sampling and Return-Based Prioritized Sampling
 Figure 4: Traffic networks used in the experiments: (a) Jinan, (b) Hangzhou, and (c) Manhattan, illustrating the varying scales and complexities of the test scenarios.
-OffLight synergistically combines importance sampling (IS) and Return-Based Prioritized Sampling (RBPS) to address both distributional shift and sample efficiency in offline MARL for TSC.
-This combined weighting scheme ensures that the learning algorithm emphasizes transitions that are both aligned with the target policy and derived from high-reward episodes, thereby enhancing the effectiveness and robustness of policy learning.
+OffLight integrates IS and RBPS to improve sample efficiency and mitigate distributional shift in offline MARL for TSC.
+IS ensures transitions align with the target policy, while RBPS prioritizes high-reward episodes, refining learning by emphasizing both relevance and quality [^4] Multiplying $w_{\text{IS},t}^{k}$ and $w_{\text{RBPS}}^{k}$ ensures that only transitions that are both relevant to the target policy and come from successful episodes are given significant weight.
+This dual prioritization enhances the effectiveness and robustness of policy learning..
 $w_{\text{combined},t}^{k}=w_{\text{IS},t}^{k}\times w_{\text{RBPS}}^{k}$ (5)
-importance sampling adjusts the weights based on policy alignment, while RBPS prioritizes transitions from successful episodes, ensuring that the model learns from both relevant and high-quality data.
-[^3] Multiplying $w_{\text{IS},t}^{k}$ and $w_{\text{RBPS}}^{k}$ ensures that only transitions that are both relevant to the target policy and come from successful episodes are given significant weight.
-This dual prioritization enhances the effectiveness and robustness of policy learning.
-Normalization: After combining, the weights are normalized within each minibatch to maintain their probabilistic interpretation and ensure numerical stability.
-$\tilde{w}_{\text{combined},t}^{k}=\frac{w_{\text{combined},t}^{k}}{\sum_{j}w_{\text{combined},t}^{j}}$ (7)
-Clipping: To prevent extremely large weights from destabilizing the training process, the combined weights are clamped to a maximum value.
-[…]
-Integration into Loss Functions: The normalized and clamped combined weights are applied to the Temporal Difference (TD) error and the Conservative Q-Learning (CQL) loss to adjust the influence of each transition during training.
-3.6 Integration with Offline RL Algorithms
+To maintain stability, the combined weights are normalized and clipped within each minibatch.
+These adjusted weights are then incorporated into the Temporal Difference (TD) error and Conservative Q-Learning (CQL) loss functions, ensuring more effective policy learning.
+III-G Integration with Offline RL Algorithms
 OffLight can easily be integrated with existing offline RL algorithms.
 CQL and TD3+BC are chosen as baseline offline RL algorithms.
-The combined weighting scheme—comprising importance sampling (IS) weights and Return-Based Prioritized Sampling (RBPS) weights—is incorporated into both the loss functions and the sampling mechanisms of these algorithms to ensure effective bias correction and sample prioritization.
+The combined weighting scheme—comprising IS weights and RBPS weights—is incorporated into both the loss functions and the sampling mechanisms of these algorithms to ensure effective bias correction and sample prioritization.
 Loss Function Adjustment: The combined weights $\tilde{w}_{\text{combined},t}$ are used to scale the loss components, ensuring that transitions from high-reward episodes aligned with the target policy have a greater impact on policy and value updates.
@@
 This prioritization enhances the efficiency of the learning process by focusing on the most informative and effective experiences.
-By integrating these techniques, OffLight provides a robust framework for learning effective traffic control policies from static datasets, even in the presence of policy heterogeneity.
-4 Experimental Setup
-We conduct experiments to evaluate OffLight’s performance across three real-world urban traffic scenarios: Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections).
-These scenarios vary in complexity and traffic demand, allowing for a comprehensive evaluation of OffLight’s scalability and robustness.
+IV Experimental Setup
+We evaluate OffLight in three real-world urban traffic scenarios  [^5] https://github.com/traffic-signal-control/sample-code/: Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections).
+These scenarios vary in size and traffic demand, allowing for a comprehensive assessment of OffLight’s scalability and robustness.
+Traffic demand is classified as low, medium (real-world data), or high by adjusting the number of vehicles entering the network proportionally.
 The experiments are designed to address the following key questions:
@@
     How does the proportion of suboptimal policy data affect overall performance?
-4.1 Traffic Scenarios
 Average Network
@@
 Manhattan 150 250 350 196
-Table 1: Traffic demand levels and network size for the three urban traffic scenarios used in the experiments, including Jinan, Hangzhou, and Manhattan.
-The experiments include three distinct traffic networks: Jinan, a small network with moderate traffic complexity (12 intersections); Hangzhou, a medium network with higher traffic flow (16 intersections); and Manhattan, a large urban network (196 intersections) with complex traffic patterns.
-Each scenario captures different levels of complexity, simulating low, medium, and high traffic demand conditions as shown in Table 1.
-4.2 Data Collection
-The dataset used for training includes traffic scenarios generated using a variety of controllers, including traditional rule-based methods (Fixed Time, Greedy, Max Pressure, Self-Organizing Traffic Light (SOTL)) and expert reinforcement learning controllers (MAPPO).
-Random policy controllers are also used to simulate suboptimal behavior.
-Each simulation runs for 1 hour, divided into 10 episodes of 6 minutes each, across three traffic demand levels (low, medium, and high).
-This results in 100 hours of simulation per controller for each demand level.
-4.3 Baselines
-OffLight’s performance is compared against several baseline offline reinforcement learning methods: Behavior Cloning (BC), Conservative Q-Learning (CQL), and TD3+BC.
+TABLE I: Traffic demand levels and network sizes for Jinan, Hangzhou, and Manhattan. Demand represents vehicles generated per incoming lane per hour.
+Traffic Scenarios: The selected networks represent different levels of complexity: Jinan (moderate, 12 intersections), Hangzhou (higher traffic, 16 intersections), and Manhattan (large-scale, 196 intersections).
+Each scenario simulates low, medium, and high demand conditions (Table I).
+Data Collection: Training data includes traffic simulations using multiple controllers: rule-based methods (Fixed Time, Greedy, Max Pressure, SOTL), expert RL controllers (MAPPO), and random policies for suboptimal behavior.
+Each simulation runs for 1 hour, divided into 10 episodes of 6 minutes each, across three demand levels, totaling 100 simulation hours per controller per demand level.
+Baselines: OffLight is compared against Behavior Cloning (BC), Conservative Q-Learning (CQL), and TD3+BC.
 Each algorithm is trained for 20k timesteps.
-4.4 Evaluation Metrics
-Performance is evaluated using two key metrics: Average Travel Time (ATT) and Queue Length (QL).
-These metrics are calculated across different traffic demand levels to assess the effectiveness of OffLight in managing traffic flow and reducing congestion.
-5 Results and Discussion
+Evaluation Metrics: Performance is measured using Average Travel Time (ATT) and Queue Length (QL), assessing traffic flow and congestion reduction across different demand levels.
+V Results and Discussion
 This section presents the evaluation of OffLight across various traffic scenarios.
 We measure performance in terms of Average Travel Time (ATT) and Queue Length (QL), highlighting the effectiveness of OffLight in learning from mixed-policy datasets.
-5.1 Performance on Mixed-Quality Data
+V-A Performance on Mixed-Quality Data
 We assess OffLight’s performance on datasets with varying proportions of expert and random policy data.
 The results are shown for three traffic networks: Jinan, Hangzhou, and Manhattan, under low, medium, and high traffic demand.
+Results show that OffLight consistently outperforms baselines, particularly in high-demand conditions, demonstrating its ability to manage congestion in large-scale networks.
 Category Algorithm Jinan Hangzhou Manhattan
@@
 OffLight (CQL) 303.24 307.34 324.00 325.27 304.55 308.56 439.82 456.10 483.78
-Table 2: Average Travel Time (ATT): Mixed Policy Dataset
+TABLE II: Average Travel Time (ATT): Mixed Policy Dataset
 Category Algorithm Jinan Hangzhou Manhattan
@@
 OffLight (CQL) 314.86 402.17 553.09 189.50 219.46 240.54 798.09 979.13 999.89
-Table 3: Queue Length (QL): Mixed Policy Dataset
+TABLE III: Queue Length (QL): Mixed Policy Dataset
 Figure 5: Performance Comparison of OffLight (CQL) and OffLight (TD3+BC) on Mixed Data showing the improvements in average travel time (ATT) across different traffic demand levels
-5.1.1 Average Travel Time (ATT)
-Figure 5 compares the performance of OffLight with baseline offline RL methods across three traffic scenarios—Jinan, Hangzhou, and Manhattan—at low, medium, and high traffic demand levels.
-Lower ATT values indicate more efficient traffic flow.
-Across all three scenarios (Jinan, Hangzhou, Manhattan), OffLight consistently outperforms both TD3+BC and CQL baselines.
-In particular, OffLight (TD3+BC) shows the best performance under high traffic demand conditions, particularly in the Jinan and Manhattan scenarios.
-In the Manhattan scenario, where traffic complexity is the highest, OffLight achieved up to a 6.9% reduction in ATT at high demand levels when integrated with CQL, and a 7.8% reduction when integrated with TD3+BC, as compared to their respective baselines.
-This highlights OffLight’s ability to handle large-scale networks effectively under congested conditions.
-In low traffic demand scenarios, the performance improvements were less pronounced, which could indicate that under lighter traffic, the margin for improvement is naturally smaller.
-The results suggest that OffLight’s incorporation of IS and RBPS allows it to prioritize transitions that are most relevant to traffic efficiency, even in datasets containing a mix of expert and suboptimal policies.
-By focusing learning on higher-quality data, OffLight is able to reduce congestion more effectively, particularly in high-demand, complex networks like Manhattan.
-5.1.2 Queue Length (QL)
+V-A1 Average Travel Time (ATT)
+Figure 5 compares OffLight to offline RL baselines.
+OffLight achieves lower ATT across all scenarios, with the greatest improvements in high-demand conditions.
+In Manhattan, OffLight reduces ATT by up to 7.8% (TD3+BC variant) and 6.9% (CQL variant) compared to their respective baselines, proving its scalability in complex networks.
+Performance gains are smaller in low-traffic scenarios, where traffic flow is already manageable.
+These results indicate that OffLight effectively prioritizes high-quality data, improving traffic efficiency in congested environments.
+V-A2 Queue Length (QL)
 Figure 6: Learning curves for the Jinan scenario under medium traffic demand, showing the convergence of queue length for OffLight and baseline methods.
-Table 2 presents the queue length results for the three traffic scenarios, indicating the efficiency of OffLight in reducing the number of vehicles queued at intersections.
-OffLight (TD3+BC) shows significant reductions in queue lengths, particularly in medium to high traffic demand scenarios.
-For example, in the Hangzhou scenario with medium traffic, OffLight (TD3+BC) reduced the queue length by 11.2% compared to TD3+BC alone.
-In Manhattan, OffLight also showed marked reductions in queue length, outperforming both baselines across all traffic levels.
-For instance, OffLight (CQL) achieved a 6% reduction in queue length under high traffic demand, which is critical in reducing congestion in highly complex urban networks.
-However, under low traffic demand, queue length reductions are less pronounced, likely because fewer vehicles are on the road, making it easier for any control policy to manage traffic effectively.
-The strong performance in queue length reduction, particularly in medium and high traffic scenarios, highlights OffLight’s ability to coordinate intersections more effectively.
-The IS mechanism ensures that OffLight learns to prioritize actions that lead to better overall traffic management, minimizing stoppages at intersections.
-Figure 6 provides a detailed visualization of the learning curves for various algorithms in the Jinan scenario on medium traffic.
-TD3+BC and CQL improve queue length reduction over time, with TD3+BC converging faster than CQL, but they still result in higher overall queue lengths compared to OffLight.
-OffLight (TD3+BC) and OffLight (CQL) both significantly outperform the other baselines, maintaining the lowest queue lengths throughout training, with OffLight (TD3+BC) performing the best overall.
-BC, TD3+BC, and CQL, on the other hand, show more variance and slower convergence.
-This indicates that OffLight is better at focusing on transitions that matter, effectively reducing congestion more consistently.
-Importance sampling (IS) ensures that transitions generated under different behavior policies are re-weighted based on their relevance to the current policy, correcting for any mismatch between the offline data and the target policy.
-This adjustment allows OffLight to more effectively utilize the dataset and focus learning on actions that lead to better traffic flow, resulting in faster convergence and significantly lower queue lengths.
-5.1.3 Behavior Representation
+Table III and Figure 6 show that OffLight significantly reduces queue lengths, especially in medium and high traffic conditions.
+In Hangzhou, OffLight (TD3+BC) decreases queue lengths by 11.2% compared to TD3+BC alone.
+In Manhattan, OffLight (CQL) achieves a 6% reduction under high traffic demand, demonstrating its capability to optimize traffic flow at intersections.
+Improvements are smaller in low-traffic scenarios where congestion is naturally less severe.
+Figure 6 highlights the learning curves for the Jinan scenario under medium traffic demand.
+OffLight (TD3+BC) and OffLight (CQL) achieve faster convergence and lower queue lengths compared to baselines, with OffLight (TD3+BC) performing best.
+TD3+BC and CQL improve over time but maintain higher queue lengths than OffLight, while BC and other baselines show greater variance and slower convergence.
+OffLight’s IS mechanism ensures that transitions are weighted based on their relevance to the learned policy, reducing reliance on low-quality data.
+This re-weighting helps OffLight make better traffic control decisions, resulting in smoother traffic flow, quicker convergence, and significantly lower queue lengths across all demand levels.
+V-A3 Behavior Representation
 Figure 7: t-SNE Representation of Actual Data and Learned Trajectories for Hangzhou scenario
-Figure 7 visualizes the t-SNE embeddings of the state-action-next state pairs and the learned latent trajectories for the Hangzhou scenario.
-The t-SNE visualizations reveal clear separation between the controllers in the latent space.
-This separation is especially evident between expert, random, and rule-based policies, with distinct clusters forming for each type of policy.
-This demonstrates the effectiveness of Graph-GMVAE in capturing the latent structure of the dataset.
-The results also indicate that OffLight is robust to variations in the quality of the dataset.
-As the proportion of expert data decreases, OffLight continues to outperform baseline methods.
-This robustness is critical in real-world settings where data quality may vary significantly across different intersections and traffic conditions.
-5.2 Ablation Studies
-5.2.1 Performance Comparison with Different Levels of Mixing
+Figure 7 visualizes t-SNE embeddings of learned latent space in the Hangzhou scenario, illustrating OffLight’s ability to differentiate between expert, random, and rule-based policies.
+The distinct clusters confirm that GMM-VGAE effectively captures structural differences in control policies, enabling more robust policy learning.
+OffLight remains effective even as the proportion of expert data decreases, consistently outperforming baselines.
+This robustness is crucial in real-world settings where dataset quality varies across intersections and traffic conditions, ensuring reliable performance despite heterogeneous data sources.
+V-B Ablation Studies
+V-B1 Performance Comparison with Different Levels of Mixing
 Figure 8: Performance Comparison of OffLight with Expert and Random Data Mix
@@
 This indicates that OffLight’s IS and RBPS mechanisms are effective in mitigating the negative effects of suboptimal data.
-5.2.2 Effectiveness of Improved Sampling Strategies
+V-B2 Effectiveness of Improved Sampling Strategies
 Figure 9: Ablation Study on CQL (top) and TD3+BC (bottom) with IS and RBPS
-We further investigate the contribution of the two components of OffLight: Importance Sampling (IS) and Prioritized Sampling (PS).
-Both strategies aim to improve learning from offline datasets by addressing different challenges: IS adjusts for the distributional shift between the behavior policy and the target policy, while RBPS focuses the learning on episodes with higher total rewards.
-IS consistently outperforms RBPS in all scenarios (Jinan, Hangzhou, Manhattan) and across all demand levels (low, medium, high).
-The gap between IS and RBPS widens as the traffic demand increases.
-This is particularly evident in the high traffic demand scenario, where IS achieves up to a 10% improvement in ATT for CQL and 9% improvement for TD3+BC in Manhattan.
-Medium and high traffic demand scenarios benefit the most from IS, as seen in Hangzhou and Manhattan, where IS outperforms RBPS by several percentage points, especially as congestion increases.
-RBPS improves ATT modestly in all scenarios, but its performance is generally lower than IS.
-In particular, in medium and high traffic demand scenarios, RBPS provides only incremental improvements (around 4-6% in ATT for both CQL and TD3+BC).
-RBPS performs more consistently across traffic demand levels, but its gains are notably smaller compared to IS, especially in more complex traffic networks like Manhattan.
-IS corrects for the distributional shift between the offline dataset (behavior policy) and the target policy, ensuring that transitions more aligned with the target policy are given higher weight.
-This correction is particularly crucial in high-demand scenarios, where suboptimal data can significantly distort learning.
-By focusing on more relevant data, IS ensures the model can make more effective decisions, leading to greater improvements in traffic flow, especially in larger, more complex networks like Manhattan.
-RBPS, while effective in prioritizing episodes with higher rewards, is more sensitive to noise in mixed-quality datasets, especially when suboptimal policies dominate certain episodes.
-This can lead to smaller performance gains compared to IS, which dynamically adjusts the importance of transitions based on their relevance to the target policy.
-5.3 Discussion
-In this study, we proposed OffLight, an offline multi-agent reinforcement learning (MARL) framework tailored for traffic signal control (TSC) using heterogeneous, pre-collected datasets.
-The experiments illustrate that OffLight not only outperforms existing baseline methods but also remains robust across different traffic scenarios and dataset compositions.
-Key insights from the results include:
-Performance with Mixed-Quality Data: OffLight consistently demonstrates superior performance in scenarios involving heterogeneous behavior policies.
-This is critical for real-world deployments where datasets often comprise both expert and suboptimal policy data.
-The Graph-GMVAE was crucial in capturing the latent behavior distributions, enabling more accurate policy learning.
-Effectiveness of Sampling Strategies: Both importance sampling (IS) and return-based prioritized sampling (RBPS) played pivotal roles in enhancing learning efficiency.
-IS proved more impactful overall, particularly in high traffic demand scenarios, by correcting for distributional shifts and focusing on transitions more aligned with the target policy.
-In contrast, RBPS provided more consistent, albeit smaller, gains by prioritizing high-reward episodes.
-Together, these strategies enabled OffLight to maximize performance, especially in complex environments like Manhattan.
-Scalability and Generalization: OffLight scaled well across different network sizes and traffic conditions, from small networks like Jinan to highly complex ones like Manhattan.
-This demonstrates the framework’s generalizability, making it suitable for deployment in diverse urban settings.
-5.3.1 Limitations:
-Training Overhead for Graph-GMVAE: Training the Graph-GMVAE, which models the behavior policies, can be computationally expensive, especially in large-scale networks with numerous intersections.
-This adds to the overall training time and resource requirements, which could limit its real-time deployment capabilities in certain urban environments.
-Dependence on Dataset Quality: OffLight’s performance is highly dependent on the quality and diversity of the offline dataset.
-If the dataset lacks sufficient expert data or is overly biased toward suboptimal policies, OffLight’s learning process may be negatively impacted.
-6 Conclusion
+Figure 9 analyzes the impact of IS and RBPS on OffLight’s performance.
+IS improves performance more than RBPS across all scenarios and traffic demand levels, with the performance gap widening as congestion increases.
+In Manhattan’s high-demand setting, IS improves ATT by up to 10% for CQL and 9% for TD3+BC, demonstrating its effectiveness in mitigating distributional shift.
+RBPS provides smaller, incremental gains (4-6% ATT reduction), especially in medium and high-demand settings.
+While RBPS prioritizes high-reward episodes, it is more sensitive to noisy data and less effective than IS in complex networks like Manhattan.
+By re-weighting transitions based on their relevance to the target policy, IS ensures that OffLight learns from high-quality data, improving traffic flow decisions in congested scenarios.
+RBPS, while useful, has a smaller impact when suboptimal policies dominate certain episodes.
+The combined approach enhances OffLight’s ability to learn from offline datasets efficiently, particularly in large-scale networks.
+V-C Discussion
+This study introduced OffLight, an offline MARL framework for traffic signal control (TSC) using heterogeneous datasets.
+OffLight consistently outperforms baselines, demonstrating robustness across different traffic conditions and dataset compositions.
+By leveraging GMM-VGAE for behavior modeling and incorporating IS and RBPS, OffLight effectively handles mixed-quality data, prioritizing high-reward transitions and mitigating distributional shifts.
+OffLight scales well across diverse network sizes, from small to large-scale urban environments.
+However, its reliance on GMM-VGAE introduces computational overhead, which may limit real-time deployment.
+Additionally, its performance depends on dataset quality, requiring a balanced mix of expert and suboptimal policies for optimal learning.
+Beyond TSC, OffLight’s framework is applicable to smart grids [25] and supply chain management [26], where offline MARL can optimize decision-making in networked systems.
+Its IS-based policy correction and graph-based modeling make it well-suited for improving coordination in distributed environments.
+VI Conclusion
 OffLight demonstrates the potential to significantly enhance traffic signal control (TSC) in urban networks through offline multi-agent reinforcement learning (MARL).
 By addressing the key challenges of heterogeneous behavior policies and distributional shifts, OffLight offers a robust solution capable of learning effective policies from mixed-quality datasets.
-The framework’s incorporation of Graph-GMVAE for behavior policy modeling, alongside importance sampling (IS) and return-based prioritized sampling (RBPS), leads to substantial improvements in traffic efficiency.
+The framework’s incorporation of GMM-VGAE for behavior policy modeling, alongside IS and RBPS, leads to substantial improvements in traffic efficiency.
 These advancements reflect the ability of OffLight to scale effectively across different network sizes and traffic conditions, making it a practical tool for offline reinforcement learning for real-world traffic management.
-Appendix A OffLight Implementation Details
-A.1 Rationale for Averaging Importance Sampling Weights
-In multi-agent settings, computing the Importance Sampling (IS) weight as a product across agents can lead to exponential scaling, causing numerical instability and disproportionately amplifying certain transitions.
-Specifically, the original formulation is given by:
-$w_{\text{IS},t}^{i}$ $=\frac{\pi_{\theta}^{i}(A_{t}\mid S_{t})}{\pi_{b}^{i}(A_{t}\mid S_{t})}$ (10)
-$=\prod_{i=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid s_{t})}$ (11)
-$\approx\prod_{i=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (12)
-To mitigate these issues, we adopt an averaging approach:
-$w_{\text{IS},t}^{i}$ $\approx\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$ (13)
-Averaging ensures that the combined IS weight remains within a manageable range, preventing the exponential growth associated with multiplication.
-This approach maintains numerical stability and ensures that each agent’s policy discrepancy contributes proportionally to the overall weight, thereby promoting balanced and stable training dynamics.
-Empirical results confirm that averaging leads to more stable convergence and prevents any single agent from disproportionately influencing the learning process.
-A.2 Pseudocode for Training Behavior Policies
-Algorithm 1 summarizes the training procedure of OffLight, highlighting the key steps where Importance Sampling and Return-Based Prioritized Sampling are applied.
-Algorithm 1 OffLight Training Procedure
-0:  Dataset $\mathcal{D}=\{\tau_{i}\}_{i=1}^{M}$, Target policy parameters $\theta$, Mixer parameters $\phi$
-1:  Train GMM-VGAE on $\mathcal{D}$ to model behavior policies
-2:  Compute total returns $G_{i}=\sum_{t=0}^{T}r_{t}$ for each trajectory $\tau$
-3:  Normalize returns $G_{i}$ to derive RBPS weights $w_{\text{RBPS}}^{i}$
-4:  Store estimated behavior policies $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})$ for each episode $i$ in $\mathcal{D}$
-5:  Store RBPS weights $w_{\text{RBPS}}^{i}$ for each episode $i$ in $\mathcal{D}$
-6:  Initialize target policy parameters $\theta_{\text{target}}\leftarrow\theta$
-7:  for each training iteration do
-8:     Sample a minibatch of transitions from $\mathcal{D}$ according to RBPS weights $w_{\text{RBPS}}^{i}$
-9:     for each transition $(\mathcal{O}_{t}^{i},\mathcal{A}_{t}^{i},\mathcal{R}_{t},\mathcal{O}_{t+1}^{i})$ in the minibatch do
-10:        Retrieve stored behavior policy probabilities $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})$ from $\mathcal{D}$
-11:        Compute IS weights $w_{\text{IS},t}^{i}$ using:
-12:         $w_{\text{IS},t}^{i}=\frac{1}{N}\sum_{n=1}^{N}\frac{\pi_{\theta}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i})}{\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})}$
-13:     end for
-14:     Combine weights:
-15:      $w_{\text{combined},t}^{i}=w_{\text{IS},t}^{i}\times w_{\text{RBPS}}^{i}$
-16:     Normalize $w_{\text{combined},t}^{i}$ across the minibatch:
-17:      $\tilde{w}_{\text{combined},t}^{i}=\frac{w_{\text{combined},t}^{i}}{\sum_{j}w_{\text{combined},t}^{j}}$
-18:     Clamp $\tilde{w}_{\text{combined},t}^{i}$ to prevent extreme values:
-19:      $\tilde{w}_{\text{combined},t}^{i}=\text{clamp}(\tilde{w}_{\text{combined},t}^{i},\text{max}=10.0)$
-20:     Compute loss $\mathcal{L}$ using:
-21:      $\mathcal{L}=\mathbb{E}_{(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\sim\mathcal{D}}\left[\tilde{w}_{\text{combined},t}^{i}\cdot\ell(\mathcal{O},\mathcal{A},\mathcal{O}^{\prime},\mathcal{R})\right]$
-22:     Update policy parameters $\theta$ using gradient descent on $\mathcal{L}$
-23:     Update target policy parameters $\theta_{\text{target}}$ using Polyak averaging or hard updates
-24:  end for
-A.3 GMM-VGAE Architecture
-A.3.1 Encoder: Capturing Spatial and Temporal Dependencies
-The encoder processes trajectory data by embedding local observations and actions through linear layers.
-GATs capture the spatial dependencies among the $N$ traffic intersections by leveraging the graph structure of the traffic network.
-Subsequently, RNNs, specifically Long Short-Term Memory (LSTM) networks, encode the temporal sequence of these embeddings to capture the action-observation history over time.
-This dual encoding ensures that both spatial relationships and temporal dynamics are effectively represented in the encoded features.
-Formally, the encoding process is defined as: {dmath} μ_t, logσ_t^2 = Encoder(O_t, A_t-1, A), z_t = μ_t + σ_tϵ,  ϵN(0, I), where:
-  • $\mathcal{O}$ represents the observations.
-  • $\mathcal{A}$ denotes the actions.
-  • $\text{Encoder}(\cdot)$ maps the input observations and actions to the parameters of the posterior distribution over the latent variables, yielding the mean $\boldsymbol{\mu}_{t}$ and log-variance $\log\boldsymbol{\sigma_{t}}^{2}$.
-  • $\mathbf{z}_{t}$ is the sampled latent variable obtained via the reparameterization trick.
-A.3.2 Latent Space: Modeling Behavior Policies
-The latent space is structured to represent the distribution of underlying behavior policies using a Gaussian Mixture Model (GMM).
-By modeling the latent variables $\mathbf{z}$ as a mixture of $K$ Gaussian components, OffLight captures the multimodal nature of behavior policies present in the heterogeneous dataset.
-This mixture model facilitates the disentanglement of different policy behaviors, enabling accurate estimation and differentiation of effective and suboptimal policies within the data.
-The GMM is defined as:
-$p(\mathbf{z})$ $=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{z}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}),$ (14)
-$q(\mathbf{z}\mid\tau_{i})$ $=\sum_{k=1}^{K}\gamma_{ik}\mathcal{N}(\mathbf{z}\mid\boldsymbol{\mu}_{ik},\boldsymbol{\Sigma}_{ik}),$ (15)
-where:
-  • $K$ is the number of mixture components.
-  • $\pi_{k}$ are the mixture weights with $\sum_{k=1}^{K}\pi_{k}=1$.
-  • $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ are the mean and covariance of the $k$-th Gaussian component.
-  • $\gamma_{ik}$ represents the responsibility of the $k$-th component for trajectory $\tau_{i}$.
-Figure 10: Evaluation of different latent dimensions for modeling behavior policy distribution on Hangzhou scenario
-Number of Latent Variables:
-The number of latent variables in the Graph-GMM-VGAE model was determined through a validation study using the Hangzhou traffic scenario.
-We tested several configurations, ranging from 2 to 32 latent dimensions, and measured the reconstruction accuracy of the behavior policies on the validation set.
-As the number of latent variables increased, the model’s ability to capture heterogeneity in the policy distributions improved.
-However, beyond 8 latent dimensions, the validation accuracy plateaued, indicating diminishing returns in performance improvement.
-Additionally, higher-dimensional models risk overfitting, particularly with smaller datasets.
-Therefore, we selected 8 latent variables as the optimal configuration, offering a good trade-off between accuracy and model complexity.
-A.3.3 Decoder: Reconstructing Policy Distributions
-The decoder reconstructs the action distributions from the latent variables by concatenating the sampled latent vectors with the encoded observations and passing them through additional GAT layers.
-This process effectively decodes the policy distributions corresponding to different latent behaviors.
-The reconstruction ensures that the latent representations accurately capture the decision-making processes of the behavior policies, enabling reliable policy reconstruction and subsequent importance sampling.
-The decoding process is formalized as:
-$\hat{\mathcal{A}_{t}}$ $=\text{Decoder}(\mathbf{z}_{t},\mathcal{O},\mathcal{A_{t-1}},\mathbf{A}),$ (16)
-$p(\mathcal{A}_{t}\mid\mathbf{z}_{t},\mathcal{O}_{t})$ $=\prod_{i=1}^{N}p(a_{t}^{i}\mid\mathbf{z},\mathcal{O}_{t}),$ (17)
-where:
-  • $\hat{\mathcal{A}}$ represents the reconstructed actions.
-  • $\text{Decoder}(\cdot)$ maps the latent variable $\mathbf{z}$ and observations $\mathcal{O}$ back to the action space.
-  • $p(a_{t}^{i}\mid\mathbf{z}_{t},\mathcal{O}_{t})$ is modeled as a categorical distribution over possible actions for agent $n$.
-OffLight employs a self-supervised learning approach, eliminating the need for explicit labels regarding behavior policies.
-By minimizing the reconstruction loss alongside the Kullback-Leibler (KL) divergence between the posterior and prior distributions in the latent space, the model learns to encode and decode behavior policies effectively.
-This self-supervised mechanism ensures that the latent space is structured to accurately reflect the diversity of behavior policies in the dataset.
-A.4 Offline RL Framework Details
-A.5 Hyperparameters
-Component Parameter Value
-Encoder LSTM hidden size 128
-FC hidden size 128
-GAT Layers 3
-Attention Heads 4
-Latent variable dimensions 8
-Hidden layer activation function LeakyReLU
-Decoder GAT Layers 3
-Attention Heads 4
-FC hidden size 128
-Hidden layer activation function LeakyReLU
-Final layer activation function Softmax
-Training Epochs 100
-Learning rate 3e-3
-Optimizer Adam
-Batch size 128
-KL divergence weight 1e-4
-Table 4: GMM-VGAE Hyperparameters and Architecture Details
-Component Parameter Value
-Agent Network RNN hidden size 128
-FC hidden size 128
-GAT Layers 1
-Attention Heads 4
-Hidden layer activation function ReLU
-CQL $\alpha$ 1.0
-Target network update interval 200 (episodes)
-TD3+BC $\alpha$ 1.0
-Target network update $\tau$ 0.005
-Training Optimizer Adam
-Learning rate 0.001
-Table 5: Offline RL Hyperparameters and Architecture Details
-We used PyTSC library [28] and the CityFlow simulator [29] to generate traffic scenarios for our experiments, while the OffLight algorithm was implemented using the OffPyMARL framework [30].
-Each offline reinforcement learning algorithm was trained for 100 epochs on a system equipped with an Intel(R) Core(TM) i9-9900K CPU @ 3.60GHz and an NVIDIA GeForce RTX 2080 GPU.
-The entire training process for OffLight took approximately 4 hours per run, with the GMM-VGAE component requiring between 5 to 20 minutes to complete its training.
-Appendix B Theoretical Analysis
-We analyze the properties of the importance sampling integration.
-This analysis underscores the robustness and effectiveness of OffLight in offline multi-agent reinforcement learning (MARL) settings, particularly for traffic signal control (TSC) applications.
-B.1 Importance Sampling Weight Consistency
-In OffLight, importance sampling is integrated to correct for the distributional shift between the behavior policy $\pi_{b}$ and the target policy $\pi_{\theta}$.
+APPENDIX
+VI-A Importance Sampling Weight Consistency
+In OffLight, IS is integrated to correct for the distributional shift between the behavior policy $\pi_{b}$ and the target policy $\pi_{\theta}$.
 We aim to show that the importance sampling weights computed using the estimated behavior policy $\hat{\pi}_{b}$ provide consistent and unbiased estimates of expectations under the target policy.
@@
 $=\mathbb{E}_{\tau\sim\pi_{\theta}}[f(\tau)].$ (12)
-This demonstrates that OffLight’s importance sampling weights yield consistent estimates when the behavior policy is accurately estimated.
-B.2 Variance Reduction via Accurate Behavior Policy Estimation
-While importance sampling provides unbiased estimates, the variance of the estimator can be high, particularly when the target and behavior policies differ significantly.
+This demonstrates that OffLight’s IS weights yield consistent estimates when the behavior policy is accurately estimated.
+VI-B Variance Reduction via Accurate Behavior Policy Estimation
+While IS provides unbiased estimates, the variance of the estimator can be high, particularly when the target and behavior policies differ significantly.
 By accurately estimating $\pi_{b}$ using the GMM-VGAE, OffLight reduces the variance of the importance weights.
-Variance Analysis: The variance of the importance sampling estimator is given by:
+The variance of the IS estimator is given by:
 $\text{Var}_{\pi_{b}}[w(\tau)f(\tau)]=\mathbb{E}_{\pi_{b}}[w^{2}(\tau)f^{2}(\tau)]-\left(\mathbb{E}_{\pi_{\theta}}[f(\tau)]\right)^{2}.$ (13)
@@
 Specifically, when $\pi_{\theta}$ is close to $\pi_{b}$, $w(\tau)$ is close to 1, minimizing the variance.
-Theorem: If $\hat{\pi}_{b}$ is a consistent estimator of $\pi_{b}$, then as the amount of data increases, the variance of the importance sampling estimator decreases.
+Theorem: If $\hat{\pi}_{b}$ is a consistent estimator of $\pi_{b}$, then as the amount of data increases, the variance of the IS estimator decreases.
 Proof: As $\hat{\pi}_{b}$ converges to $\pi_{b}$, the distribution of $w(\tau)$ becomes tighter around 1, reducing the variance.
 This is formalized by the Law of Large Numbers and convergence properties of consistent estimators.
-B.3 Computational Complexity and Scalability
-OffLight leverages Graph Neural Networks (GNNs) to capture spatial dependencies among agents. The computational complexity per forward pass is:
-$\mathcal{O}(L\cdot(N+E)\cdot d),$ (25)
-where: - $L$ is the number of GNN layers.
-- $N$ is the number of nodes (agents).
-- $E$ is the number of edges (inter-agent connections).
-- $d$ is the dimensionality of the embeddings.
-In sparse traffic networks, $E=\mathcal{O}(N)$, so the complexity scales linearly with $N$.
-This demonstrates that OffLight is scalable to large networks, such as the Manhattan scenario with 196 intersections.